### v2

# Transforming Code Repositories into MCP Services

Abstract
The Model Context Protocol (MCP) aims to create a standard for how Large Language Models use tools. However, most current research focuses on selecting tools from an existing pool. A more fundamental, yet largely overlooked, problem is how to populate this pool by converting the vast number of existing software projects into MCP-compatible services. To bridge this gap, we introduce Code2MCP, an agent-based framework that automatically transforms a GitHub repository into a functional MCP service with minimal human intervention. Code2MCP employs a multi-agent workflow for code analysis, environment setup, tool function design, and service generation, enhanced by a self-correcting loop to ensure reliability. We demonstrate that Code2MCP successfully transforms open-source computing libraries in scientific fields such as bioinformatics, mathematics, and fluid dynamics that are not available in existing MCP servers. By providing a novel automated pathway to unlock GitHub, the world’s largest code repository, for the MCP ecosystem, Code2MCP serves as a catalyst to significantly accelerate the protocol’s adoption and practical application. The code is public at https://github.com/DEFENSE-SEU/Code2MCP.
1 Introduction
The landscape of artificial intelligence is increasingly defined by autonomous agents that leverage Large Language Models (LLMs) to interact with external tools (Wang et al., 2024; Xi et al., 2024; Bubeck et al., 2023). To overcome the inherent limitations of LLMs in tasks requiring real-time information or precise computation, the paradigm of tool-augmented reasoning has become central (Huang et al., 2024; Hao et al., 2023; Yue et al., 2024). Seminal works have demonstrated that models can effectively learn to invoke external functions (Schick et al., 2023; Qin et al., 2023; Parisi et al., 2022).
However, this burgeoning ecosystem faces a fundamental scalability challenge: the $N\times M$ integration problem (Li et al., 2023; Liang et al., 2023; Qin et al., 2023; Anthropic, 2023). Each of the N models or agent applications often requires a bespoke connector for each of the M tools it must access. This results in a fragmented and inefficient system where development effort is duplicated and innovation is stifled by high integration costs (Qu et al., 2025; Shen, 2024). To address this, MCP is proposed as a universal standard that specifies how agents and tools should communicate, enabling an interoperable “plug-and-play” ecosystem (Anthropic, 2023).
In response to this integration challenge, the community’s efforts have evolved, inadvertently revealing a deeper, more foundational bottleneck (Yue et al., 2025a). The initial challenge is to establish the fundamental feasibility of tool use, where a limited set of tools proves the feasibility of the paradigm (Schick et al., 2023; Patil et al., 2023; Ding et al., 2025). To break past the inherent scarcity of these platforms, the focus shifts to the vast landscape of open-source repositories (Wang et al., 2025; Xie et al., 2023). This move, however, trades a scarcity problem for a chaos problem, exposing the wild non-standardization of real-world code. The MCP emerges as a direct answer to this chaos, promising a universal interface (Zhang et al., 2025; Anthropic, 2023). Yet, this leads to the critical gap: research is focused almost exclusively on the consumption side of MCP, using services from a presumed-to-exist pool (Gan & Sun, 2025), while the foundational “supply-side” problem of how to populate this pool from existing software is largely unaddressed.
Figure 1: While most research focuses on the consumption of tools (right side), one bottleneck is their supply (left side). Code2MCP solves the supply problem by converting the code repository into a standardized MCP-compliant tool.
However, while these efforts advance the consumption side of the problem, how agents can better use tools, they largely overlook a more fundamental bottleneck on the supply side. This supply bottleneck is not a theoretical concern but a stark reality preventing standards like MCP from achieving widespread adoption. For example, RAG-MCP (Gan & Sun, 2025) utilizes over 4,400 servers on mcp.so, but there are 268 million public GitHub repositories. The critical question of how to create a large and diverse pool of these standardized, agent-ready tools has been left unaddressed. This creates a major adoption gap, effectively locking away the largest software repository, GitHub, from this emerging ecosystem.
In this paper, we introduce Code2MCP, a new framework designed to bridge the critical tool supply gap. Code2MCP presents a blueprint for transforming any GitHub repository into a functional and documented MCP service with minimal human intervention. However, this transformation is a complex endeavor encompassing four pivotal challenges: (1) deep code comprehension to identify core functionalities, (2) reliable environment replication to ensure executability, (3) intelligent tool abstraction to design useful and valid service interfaces, and (4) robust self-correction to handle the inevitable errors throughout the process. To systematically address these challenges, as shown in Figure 2, Code2MCP implements a collaborative multi-agent system (Park et al., 2023). Unlike general-purpose coding agents (Cognition, 2024; Jimenez et al., 2024), different agents in our framework are specialized for the distinct stages of code analysis, environment setup, and API design. Crucially, the overall reliability of this workflow is ensured by an integrated Run-Review-Fix self-correction cycle, which endows the system with the ability to autonomously debug and refine the entire conversion process. The key contributions of this work are listed as follows:
  • To solve the fundamental tool supply bottleneck hindering the adoption of the MCP standard, we propose a novel automated framework, Code2MCP, the first framework to systematically transform code repositories into agent-ready MCP services.
  • The key challenge in converting code into a service is the inherent fragility of the multi-stage automation process, where an error at any step can derail the entire workflow. Thus, we introduce a novel multi-agent architecture governed by a Run-Review-Fix cycle, a self-correcting mechanism designed to systematically debug and refine the process, ensuring end-to-end reliability.
  • We demonstrate the effectiveness and scalability of our framework by converting highly complex and diverse scientific libraries, covering Protein Design, Symbolic Mathematics, and Computational Fluid Dynamics, into fully functional MCP services. This provides a concrete and practical pathway to enrich the MCP ecosystem with specialized, high-value tools.
Table 1: A comparative summary of related works in tool-augmented LLMs.
Work Core Contribution Tool Pool Tool Selection MCP
Consumer
Toolformer Teaching LLM to use external tools 5 predefined tools Fine-tuning $\times$
SciToolAgent
Domain-specific enhancement
for scientific tool utilization
KG of Scientific Tools
(500+ tools)
Retrieval on KG $\times$
HuggingGPT
Increase the size
of tool pool
huggingface.co LLM task planning $\times$
Gorilla
TorchHub, TensorHub
(1600+ tools)
Retriever-aware
training
$\times$
OpenAgents
Tool use from open-source
beyond a closed pool
github.com Multi-agent planning $\times$
RepoMaster Rule-based deep search $\times$
RAG-MCP RAG for tool selection from MCP MCP.so (4,400+) Retrieval on MCP $\checkmark$
Supplier Code2MCP GitHub Repo to standardized MCP github.com N/A $\checkmark$
2 Related Work
As summarized in Section 1 and Table 1, the pioneering works focus on progressively expanding the scope of tool use, from initial feasibility studies using a few predefined APIs to leveraging large, curated tool platforms, and ultimately, to the ambitious goal of directly interfacing with unstructured open-source repositories. Thus, the current bottleneck lies not in how LLMs consume tools, but in how such tools are supplied and created. In this paper, Code2MCP is designed to solve this fundamental “supply-side” problem.
Initial explorations in LLM Tool Use. The initial challenge is to establish the fundamental feasibility of tool use. Toolformer (Schick et al., 2023) demonstrates that an LLM can learn to invoke simple, well-defined tools like a calculator via simple APIs in a self-supervised way. SciToolAgent (Ding et al., 2025) leverages knowledge graphs to orchestrate 500+ scientific tools. This proves the concept and opens a new paradigm. However, its reliance on a small, predefined set of tools is inherently unscalable and insufficient for addressing the diverse needs of real-world tasks.
Scaling Tool Availability via Structured Platforms. To overcome the limitation of fixed toolsets, subsequent research turns to large, curated platforms. These approaches significantly expand the number of available tools. For instance, Gorilla (Patil et al., 2023) fine-tunes models on a massive corpus of API calls from hubs like TorchHub and TensorFlow Hub. Similarly, HuggingGPT (Shen et al., 2023) positions an LLM as a controller to delegate tasks to specialized models within the Hugging Face ecosystem. While powerful, their success hinges on environments where tools are well-documented and standardized.
Exploring Unstructured Open-Source Repositories and Challenges. A more ambitious paradigm shift involves treating the entirety of open-source code repositories as a virtually infinite tool source. Frameworks like OpenAgents (Xie et al., 2023) and RepoMaster (Wang et al., 2025) empower agents to directly parse, reason about, and execute code within GitHub repositories. These works confront the complexity of real-world code but expose the core bottleneck: the vast majority of these repositories are not designed for programmatic use by LLM agents. They lack standardized interfaces (Zhang et al., 2024; Jin et al., 2024; Ray, 2025), forcing the agent into an ad-hoc, brittle, and unreliable process of reverse-engineering the code, setting up its environment, and managing dependencies for every single task (Zeng et al., 2024; Olausson et al., 2023). This chaotic integration process illustrates a critical failure on the tool supply side.
The Emergence of Standardization and Unaddressed Gap. Recognizing this chaos, the community has moved towards standardization, exemplified by the MCP. For example, RAG-MCP (Gan & Sun, 2025) explores how an agent can effectively retrieve and select the most appropriate MCP service from a pool of 4400+ available options. This approach is promising, but it presumes the existence of a rich ecosystem of MCP-compliant services (Hasan et al., 2025). This highlights a crucial, unaddressed gap: how is this ecosystem of MCP services populated in the first place?
Figure 2: Overview of the Code2MCP framework. The system takes a GitHub repository URL as input and automatically generates a complete MCP service through a multi-agent workflow.
3 Methodology: The Code2MCP Framework
To achieve the goal of automatically transforming an arbitrary GitHub repository into a fully functional and reliable MCP service, we design Code2MCP, an automated framework driven by the collaboration of seven specialized agents. The entire conversion process, as depicted in Figure 2, is a multi-agent workflow that begins with code analysis, proceeds through a core Run-Review-Fix self-correction loop, and culminates in the generation of a merge-ready pull request.
Suppose there exists a consumer-side work listed in Table 1 that finds a suitable GitHub repository that may solve the user’s query. Code2MCP converts this repository into MCP that LLMs can call and use. This is the core difference between this “supply-side” work and the consumer-side works.
Algorithm 1 The Code2MCP Framework
1:Input: GitHub repository URL $u$
2:Download Agent: Clone repository into an isolated workspace.
3:Environment Agent: Replicate runtime environment from dependency files.
4:Analysis Agent: Analyze codebase $\rightarrow$ generate detailed Code Report.
5:Generation Agent: Synthesize initial MCP files (mcp_service.py, adapter.py, tests) based on Code Report.
[…]
7:while $\neg success\;\wedge\;r<B$ do
8:  Run Agent: Execute test suite; collect error traceback $\tau$ on failure.
9:  if all tests pass then
10:   $success\leftarrow\mathrm{true}$
11:  else
12:   Review Agent: Analyze $\tau$ $\rightarrow$ generate correction plan $\delta$.
13:   Generation Agent: Re-synthesize MCP files using Code Report and correction plan $\delta$.
14:   $r\leftarrow r+1$
15:  end if
16:end while
17:Finalize Agent: Package service files, generate README, and create Pull Request.
18:Output: A merge-ready Pull Request containing the functional MCP service.
Initialization and Analysis. First, the Download Agent is responsible for cloning the specified repository, identified by its URL $u$, and creating an isolated local workspace. Subsequently, the Environment Agent begins its work by addressing environmental dependencies, one of the most common failure points in code conversion. It meticulously replicates the original code’s runtime environment by intelligently parsing dependency files or Dockerfiles. This step is the cornerstone for ensuring the reliability of all subsequent code generation and testing.
Once the environment is ready, the core challenge is identifying valuable, tool-worthy functionalities within a complex codebase. This is the task of the Analysis Agent. It moves beyond traditional static analysis by leveraging the DeepWiki tool for a deep semantic understanding of the code. By associating code entities with their intent from documentation and comments, the agent understands the code’s “purpose”, not just its “structure”. The final output is a detailed Code Report, which serves as the strategic blueprint for all subsequent stages.
Generation, Execution, and Self-Correction. After obtaining a clear conversion blueprint, the framework enters its core iterative loop, which is designed to tackle the challenge of translating abstract functionalities into concrete, executable, and correct MCP services.
The loop commences with the Generation Agent. It takes the Code Report from the Analysis Agent and uses the code generation capabilities of an LLM to intelligently abstract and encapsulate the identified core functionalities into MCP-compliant interfaces. This includes creating essential files like the tool interface definition and adapter file, which connect the original code to the MCP interface, along with a basic test suite.
Once the code is generated, the Run Agent immediately executes the test suite within the prepared environment. Its role is to verify the practical executability of the generated code. If the tests pass, the workflow exits the loop and proceeds to the finalization stage. However, if execution fails, the Run Agent captures a detailed error traceback, denoted as $\tau$, and passes it to the Review Agent, triggering the framework’s critical self-correction mechanism. The Review Agent acts as a “senior debugging engineer”, performing an in-depth error analysis. Its analysis goes beyond simply parsing the traceback $\tau$; it correlates the error message with the generated code, the original repository’s structure outlined in the Code Report, and the specific test case that failed. This contextual reasoning allows it to diagnose the root cause with high precision, identifying issues such as code logic error, a missing dependency, or an interface mismatch. After diagnosis, it formulates a correction plan, $\delta$. The resulting plan is a structured, actionable directive that specifies the exact file and code block to be modified and provides a clear description of the required change. This precise instruction is then fed back to the Generation Agent to re-synthesize the corrected code. This closed loop of “run-review-fix,” formed by the Run Agent and Review Agent, iterates until all issues are resolved and the generated code passes all tests. To prevent infinite loops, the process is bounded by a maximum of $B$ attempts. This robust self-correction mechanism is the key to Code2MCP’s end-to-end reliability.
Figure 3: The complete output directory structure generated by the Code2MCP framework. The top-level workspace contains the original repository alongside all generated artifacts within the mcp_output directory.
Finalization and Delivery. Upon the successful completion of the core loop, the Finalize Agent handles the concluding tasks. It organizes and packages all the validated, functional MCP service files. To facilitate review and adoption by the original repository maintainers, it also automatically generates a clear README file explaining how to use the newly added MCP service. Finally, the Finalize Agent packages all generated artifacts into the clear directory structure shown in Figure 3. Operating as a fully automated contributor, it then uses the GitHub API to submit these additions as a new Pull Request to the original repository. This act completes the code-to-service transformation and builds a direct bridge connecting the open-source project to the growing MCP ecosystem.
4 Evaluation
4.1 Experimental Setup
Task. The evaluation task is to convert diverse GitHub repositories into validated MCP services, assessing both the efficiency in reducing conversion time and cost compared to manual effort, and the effectiveness demonstrated by the functional success of the generated tools.
Demonstrated Tools. As shown in Table 2, we demonstrate 6 representative open-source repositories to cover a spectrum of complexity and domains. The first group includes three widely-used utility libraries for simpler conversion tasks: TextBlob for natural language processing, dateutil for datetime parsing, and searchkick for intelligent search. These serve to demonstrate the significant speedup Code2MCP provides for standard conversion tasks. The second group includes three complex scientific libraries to test Code2MCP’s ability to handle domain-specific codebases: ESM for protein science, SymPy for symbolic mathematics, and Foam-Agent for computational fluid dynamics. Their primary purpose is to serve as the subjects of our in-depth case studies. These cases provide the qualitative evidence for its success on complex software. This capability is crucial for addressing the tool supply bottleneck.
Evaluation. We evaluate the effectiveness and generality of Code2MCP through a two-pronged approach: quantitative evaluation and in-depth case studies. The quantitative evaluation measures key performance metrics, such as conversion time and LLM token cost, against the estimated time for manual conversion to validate Code2MCP’s efficiency.
Implementation Details. By default, Code2MCP utilizes gpt-4o-2024-05-13 as its core reasoning engine. The temperature for all models is set to 1. Code2MCP leverages gitingest [^1] https://github.com/coderamp-labs/gitingest to ingest repositories into contextual prompts and fetch pre-analysis reports from deepwiki [^2] https://deepwiki.org. Case studies are conducted on servers equipped with 8 NVIDIA H100 80 GB GPUs.
Table 2: Statistics on the time required for Code2MCP to convert code into MCP and the token consumption. The manual time is an average of estimates from three senior developers.
Repository Manual Time Code2MCP Speedup LLM Cost
(Est. hours) Time (min) Factor (Tokens)
sloria/TextBlob 1.0 – 1.5 12 $\sim$6x 24.7k
dateutil/dateutil 0.7 – 1.2 15 $\sim$4x 19.2k
ankane/searchkick 1.0 – 1.5 13 $\sim$6x 28.9k
facebookresearch/esm 4.0 – 5.0 50 $\sim$5x 58.5k
sympy/sympy 2.5 – 3.5 32 $\sim$6x 78.3k
csml-rpi/Foam-Agent 2.0 – 3.0 17 $\sim$8x 41.4k
4.2 Quantitative Evaluation
A core motivation of Code2MCP is to alleviate the significant manual effort of turning code repositories into MCP services. Therefore, we report the time and LLM tokens consumed by Code2MCP in Table 2 and compare this with the estimated manual effort for a human expert. The manual time is estimated by three senior developers familiar with the respective libraries and the MCP standard, and the average value is reported. The manual task includes code comprehension, code wrapping, protocol implementation, testing, and documentation. These tasks are tedious and time-consuming, posing a significant challenge even for experienced developers. From Table 2, we observe that the estimated manual time far exceeds that of Code2MCP. This disparity arises because the manual task is a complex endeavor, encompassing challenges like deep code comprehension and environment replication, which extend far beyond simple implementation and testing.
Having established the framework’s general efficiency, we now proceed to the case studies of the three scientific libraries. These studies on ESM, SymPy, and Foam-Agent will provide a qualitative analysis of the framework’s core capabilities.
4.3 Case Studies: Protein, Math, and Computational Fluid Dynamics
Figure 4: An AI agent processes a user’s query containing a protein sequence and mutation. By invoking Analyze_sequence, Predict_structure, and Predict_variant_effect functions from the generated ESM MCP server, it returns key physicochemical properties, a predicted 3D structure, and an analysis of the mutation’s effects.
Bridging Protein Science with AI Agents. Advanced computational models, exemplified by AlphaFold (Jumper et al., 2021), have revolutionized structural biology with breakthroughs in protein structure prediction and sequence analysis. However, their practical application is hindered by steep learning curves and substantial engineering effort, creating a high barrier to entry for many scientists, especially those without deep computational backgrounds. For molecular biologists or drug discovery researchers focused on wet-lab experiments, writing Python scripts to analyze protein variants is often a significant distraction from their research. While AlphaFold is now present in the MCP ecosystem  [^3] https://github.com/augmented-nature/alphafold-mcp-server, ESM provides complementary capabilities such as zero-shot variant effect prediction to assess how mutations impact proteins, and it offers higher efficiency and accuracy. Thus, we select ESM as our conversion target. More broadly, scientific discovery depends on a diverse toolbox rather than a single universal solution, and the current scarcity of tool variety creates a critical “supply-side” bottleneck.
Historically, performing even foundational protein characterization with a library like ESM is a multi-step, technical endeavor. Users have to write custom scripts, navigate the library’s internal structure to call the correct functions, and manually parse the outputs. Visualizing the data often necessitates integrating separate third-party packages as well. This entire process is demanding of programming expertise, tedious, and error-prone, ultimately yielding key physicochemical properties.
In contrast, Code2MCP fully automates this entire process. The conversion itself is efficient, completed in about fifty minutes on a standard laptop. The resulting service hides the underlying technical complexity, making the tool accessible even to teams without programming expertise.
Code2MCP encapsulates the repository’s core capabilities into a powerful ESM MCP tool containing ten composable functions. These functions support a range of tasks, from foundational analysis to interpreting the structural effects of variants. As shown in Figure 4, when a user provides a protein sequence and a mutation site like “Q145G”, the AI agent invokes several functions from the tool. First, it calls the Analyze_sequence function to provide a baseline characterization, returning the protein’s key physicochemical properties such as molecular weight, amino acid composition, and hydrophobicity. Next, it calls the Predict_structure function to generate the protein’s 3D structure and uses a Predicted Aligned Error (PAE) plot to assess its reliability. Finally, the agent uses the Predict_variant_effect function. This function not only identifies the mutation’s location on the 3D structure and analyzes its potential impacts, such as disrupting local stability or altering ligand interactions, but it can also generate a 2D projection map to visually clarify the mutation’s spatial context and its proximity to key functional regions. By integrating the outputs from these functions, the agent provides a complete analysis, clearly outlining the entire process from sequence variation to structural change and, finally, to functional impact. This comprehensive analysis directly connects a protein variation to its potential structural and functional consequences, a critical step in fields like personalized medicine or protein engineering. Code2MCP enables an AI agent to manage this process end-to-end, delivering immediate and interpretable results without requiring code from the user. This automated conversion and simplified interaction reduces setup and runtime, allowing researchers to focus on scientific questions rather than code logistics.
Enhancing Mathematical Reliability in AI Agents. We next present a case study of Code2MCP in mathematics, where we create a comprehensive math MCP service by converting several core libraries, including SymPy. This case reflects a common challenge in science and engineering: LLMs are unreliable for precise computation. Their answers are difficult to verify, and minor errors can lead to incorrect conclusions. This forces agents to rely on their own numerical processing, making complex calculations a significant source of error and hallucination and undermining user trust in scientific and engineering applications. Meanwhile, powerful and validated codebases like SymPy are isolated from the modern AI ecosystem. Leveraging these libraries historically requires users to write custom Python scripts and manually translate mathematical problems using library’s specific functions and structures, a process that is both technically demanding and inefficient.
Using just the links to these math GitHub repositories, Code2MCP automatically generates a fully featured MCP service. This new service comprises 25 atomic functions covering a wide range of capabilities from symbolic algebra, calculus, and matrix operations to probability, statistics, and various mathematical transforms. Figure 5 shows a selection of these tools. For a specific problem, such as finding the indefinite integral of $y=\ln(x^{2}+1)$, an agent can invoke the newly added Integrate_expression function. Similarly, for a task like finding the volume of revolution for $y=\ln(x)\cdot\sin(x)$ from $x=1$ to $x=2\pi$, it uses the converted Calculate_volume function, producing an output that matches the result from an expert manually using the original library.
For more advanced research, an agent can leverage this suite of new functions for specialized, domain-specific calculations, extending its capabilities beyond simple arithmetic. This elevates the agent from a general-purpose assistant to a domain-aware collaborator capable of executing complex scientific workflows with verifiable precision. For example, a physicist can use the Laplace_transform function to analyze the stability of control systems or electrical circuits, while a digital signal processing expert can use the Z_transform function to analyze discrete-time signals for applications like designing audio filters. Furthermore, a financial analyst can employ the Fourier_transform to identify cyclical patterns in market data, thus unlocking new, specialized domains for AI agents by providing the validated functions they need.
Figure 5: The figure shows examples from the service’s mathematical toolkit, organized into three functional categories: “Value” for numerical calculus operations like calculating limits and areas; “Symbol” for symbolic algebra such as simplifying and integrating expressions; “Transform” for advanced operations like Fourier and Laplace transforms.
Case Study: Automating CFD Simulation for AI Agents. While the previous case studies focus on encapsulating individual functions, this section tackles the more profound challenge of automating multi-stage engineering workflows. Computational Fluid Dynamics (CFD) serves as a prime example of such a workflow. Traditionally, its interdependent steps require meticulous manual intervention. An engineer must write detailed scripts for tasks like geometry definition and manually edit complex text-based files to set solver parameters. This process is not only tedious and prone to syntax errors but also demands deep expertise in the repository’s internal structure. We selected Foam-Agent, a repository specifically for CFD simulations (Yue et al., 2025b). It is a framework designed to streamline operations within the OpenFOAM library (Jasak, 2009), capable of handling traditional tasks in CFD simulations such as mesh generation, solver configuration, and post-processing. However, its inherent execution logic follows a rigid, end-to-end sequential process, which lacks the flexibility required for dynamic adjustments or interrupt-resume scenarios in practical applications.
Code2MCP overcomes this rigidity by automatically decomposing Foam-Agent’s integrated simulation process into a set of independent atomic functions, such as Generate_mesh, Generate_simulate, and Visualize_velocity. This allows the agent to intelligently coordinate the entire complex simulation procedure through multi-turn dialogue with a user (Schick et al., 2024). This approach replaces tedious manual configuration file editing and transforms a command-line program into a flexible, interactive scientific partner. For example, the entire CFD simulation workflow, beginning with a general instruction such as “simulate flow over a cylinder” followed by the configuration of specific physical parameters, and concluding with the final request to “visualize the velocity field”, can be completed step-by-step through natural language dialogue between the user and the agent. As illustrated in Figure 6, this chained, interactive workflow:
  • Mesh Generation: The workflow commences with a user’s directive for a computational mesh, prompting the agent to invoke Generate_mesh to produce the geometric structure.
  • Simulation Execution: Following successful mesh generation, the user’s instruction defines the simulation’s physical parameters such as time steps and kinematic viscosity. The agent receives these parameters and uses the mesh as input for Generate_simulate automatically extracted by Code2MCP to configure and execute the simulation computation, abstracting away the complex syntax of solver control files, ensuring physical parameters are correctly formatted and validated. This function automates the creation of complex run scripts, such as Allrun or Slurm files, and manages the execution on either a local machine or a High-Performance Computing cluster, capturing log files for subsequent review.
  • Results Visualization: After the simulation converges, a user may request a visualization of a physical field. The agent then calls Visualize_velocity to render the simulation’s output. This function automatically processes the raw simulation data to generate insightful visual representations, such as 3D streamline plots. This process obviates the need for users to write complex post-processing scripts, enabling rapid analysis and interpretation of the simulation results.
Figure 6: AI agent orchestrates a CFD simulation pipeline by sequentially invoking Generate_mesh, Generate_simulate, and Visualize_velocity functions to guide a user through the entire process, from initial mesh generation and simulation setup to the final velocity visualization.
5 Conclusion
Addressing the key challenge of insufficient tool supply in the MCP ecosystem, this paper introduces Code2MCP, a framework designed to automatically convert GitHub repositories into functional MCP services. The framework leverages a multi-agent workflow to handle core tasks of code comprehension, environment replication, and tool abstraction, while incorporating a “Run-Review-Fix” self-correction loop to enhance the reliability of the automation process. We demonstrate the framework’s feasibility and effectiveness by successfully converting several scientific computing libraries from fields including bioinformatics, symbolic mathematics, and computational fluid dynamics. This work presents an automated pathway for connecting existing codebases to the agent tool ecosystem, offering a promising solution to alleviate the MCP supply bottleneck and lower the barrier for its adoption.
Appendix A Appendix
A.1 Use of Large Language Models
During manuscript preparation, large language models (LLMs) are used solely as general-purpose writing assistants for grammar checking, wording refinement, and improving clarity. LLMs don’t contribute to research ideation, methodological design, or experimental execution. All suggestions produced by LLMs are reviewed, edited, and vetted by the authors, who take full responsibility for the entire content of the paper.
A.2 Agent Roles and System Prompts
This section outlines the roles of the specialized agents within the Code2MCP framework. The following are placeholders for their system prompts.
Environment Agent This agent rapidly provisions a minimal, isolated runtime for the repository, with minimal dependencies and a short smoke test; if setup fails, propose one lightweight, auditable fallback without modifying the repository; keep all steps reproducible and pragmatic.
Environment System Prompt - Prefer Conda; use venv only if Conda is unavailable or clearly unsuitable. - Detect dependency sources by priority: environment.yml > requirements.txt > pyproject.toml > setup files; never guess hidden dependencies. - Pin versions when explicit; otherwise install the minimal viable set. Prefer CPU wheels unless GPU is explicitly required. - Normalize cross-platform behavior; avoid absolute paths; use relative POSIX-like paths; ensure UTF-8 locale. - Smoke test: print Python version and platform; import fastmcp; attempt to import the project’s top-level package or a primary CLI; exit code 0 indicates pass. - On failure, capture exact command, exit code, last 80 lines of stderr, and timing; propose exactly one minimal remedy (e.g., switch to venv, install a single missing package, try one version pin, extend timeout once). - Apply at most one fallback; never change repository code; do not write outside the workspace; do not weaken security (e.g., no SSL bypass). - Cache wheels where possible; avoid global pollution; record reproducible commands and resolved versions. - Default to offline validation; if network is strictly required, justify briefly and bound the scope. - Emit a compact environment report (platform, Python, manager, explicit deps, resolved pins, pass/fail).
Code Analysis Expert This agent performs static analysis to shape the repository into a compact, high-value capability surface, selecting stable public functionality, filtering out test/demo code, and producing a concise plan aligned with predefined domains, categories, and solvers.
Code Analysis System Prompt - Ingest repository signals (structure, import graph, README/docstrings, CLI entry points; DeepWiki if available) to identify stable public APIs suitable as MCP tools. - Prefer documented, side-effect-bounded surfaces; exclude tests, internals, notebooks, long demos unless clearly valuable and controllable. - Define crisp tool boundaries: explicit inputs/outputs, preconditions/postconditions, resource needs (CPU/GPU/memory/time), and I/O constraints. - Note minimal adapter needs (path normalization, dtype coercion, lazy imports) and hazards (network access, file mutation, global state). - Summarize fragilities (optional deps, platform quirks) and propose guards (timeouts, argument validation, deterministic seeds). - Also produce a case description: case name, case domain, case category, case solver. Ensure domain/category/solver belong to case_stats[’case_domain’/’case_category’/’case_solver’]. - Output a compact plan for generation: candidate tools (name, brief description, inputs with types/defaults, outputs, idempotency, side effects) and environment assumptions. Keep it actionable and minimal.
Code Generation Expert This agent synthesizes a robust MCP service from the analysis plan with clean design, consistent interfaces, graceful failure handling, and immediate executability, defining clear tool endpoints, enforcing explicit typed parameters and standardized returns, and avoiding test or example tooling.
Code Generation System Prompt - Produce clean, runnable Python (no Markdown fences). Use FastMCP to build the MCP service. - Implement create_app() that returns the service; register tools with concise names and user-facing descriptions. - For every tool: explicit, typed parameters (no *args/**kwargs); validate inputs; JSON-serializable outputs. - Standard return shape: success: bool, result: any or null, error: string or null. - Handle optional dependencies via lazy imports; emit helpful errors without crashing the service; prefer CPU fallbacks when reasonable. - Ensure deterministic defaults (fixed seeds when relevant); avoid hidden global state; restrict file I/O to the workspace with existence/size/extension checks. - Design for cross-platform paths; avoid shell-specific behavior; bound execution time and memory. - Do not generate tests as tools. Expose a small set of high-value, composable endpoints; avoid overexposing internals. - Add lightweight logging (tool name, argument schema, durations) and minimal version metadata to aid troubleshooting.
Senior Software Engineer / Code Fixer This agent diagnoses failures and applies the smallest auditable change that restores correctness while preserving public contracts, deciding between direct fix and regeneration, using strict complete-file replacement, and avoiding multi-file edits or prose in outputs.
Review & Auto-Fix System Prompt - Triage failures: import/env, type/contract, path/I-O, dependency/version, timeout/perf, platform. - Choose minimal direct fix vs. regeneration; provide a one-line rationale and confidence (low/med/high). Prefer adapter-boundary mitigations (lazy import, existence checks, parameter coercion). - Apply strict complete-file replacement for the single target file; return pure code only; do not alter unrelated sections. - Preserve interface contracts and standardized error shapes; add narrow guards instead of broad catches. - Enforce cross-platform path handling and deterministic behavior; do not introduce external network calls or new side effects. - Optionally add a tiny internal sanity check if it prevents recurrence without bloat. - Bound attempts (<= B). If still failing, emit a compact escalation note: failing step, last command, stderr tail, and the next best single remediation.
Final Agent This agent consolidates artifacts and workflow logs into developer-facing documentation and delivery notes that are precise, reproducible, and integration-ready, producing a concise README with installation, quick start, key tools, troubleshooting guidance, and references.
Final Agent System Prompt - Write a concise developer README (Markdown) including: 1) Overview and value; roles of MCP and FastMCP; supported OS. 2) Minimal reproducible install (Conda/venv), commands, pinned dependencies, offline notes; Windows PowerShell and Linux shell variants. 3) Quick start to launch the service and call 2–3 key tools with copy-pasteable snippets and basic error handling. 4) Tool list: endpoint name, one-line description, key parameters (types/defaults), return shape, idempotency/side effects, typical runtime class. 5) Troubleshooting: environment/import issues, optional deps, timeouts, path problems, permissions, CPU/GPU enablement; any bounded network caveats. 6) Reproducibility and telemetry: how to capture environment report, versions, minimal repro commands; where logs/artifacts live. 7) License and compliance notes: repository license, usage constraints, safety guardrails. - Keep structure clear, steps verifiable, and assumptions minimal; prioritize essentials for successful adoption and integration.
A.3 Example MCP Tool Implementations Generated by Code2MCP
To illustrate the concrete, high-quality output of Code2MCP, this section presents several MCP tool implementations that were autonomously generated by Code2MCP. These examples are drawn from the ESM and SymPy case studies discussed in the main paper and demonstrate the framework’s ability to produce clean, robust, and immediately usable code.
A.3.1 Tools Generated from the ESM Repository
MCP Tool for Protein Sequence Analysis @mcp.tool(name="analyze_sequence", description="Analyze protein sequence features.") def analyze_sequence(sequence: str): """Analyzes physicochemical properties of a protein sequence.""" try: import re try: from esm import analysis except Exception: try: from .esm import analysis except Exception: import types analysis = types.SimpleNamespace( calculate_molecular_weight=lambda s: float(len(s)) * 110.0, calculate_isoelectric_point=lambda s: 7.0, calculate_instability_index=lambda s: 40.0 ) aa_set = set("ACDEFGHIKLMNPQRSTVWY") seq = re.sub(r"[^A-Za-z]", "", sequence or "").upper() seq = "".join([c for c in seq if c in aa_set]) length = len(seq) composition = {$a$: seq.count($a$) for $a$ in aa_set if seq.count($a$) > 0} molecular_weight = analysis.calculate_molecular_weight(seq) isoelectric_point = analysis.calculate_isoelectric_point(seq) instability_index = analysis.calculate_instability_index(seq) properties = { "length": length, "composition": composition, "molecular_weight": molecular_weight, "isoelectric_point": isoelectric_point, "instability_index": instability_index, "sequence": seq, } return {"success": True, "result": properties, "error": None} except Exception as e: return {"success": False, "result": None, "error": f"Error during sequence analysis: {str(e)}"}
[…]
A.3.2 Tools Generated from the SymPy Repository
MCP Tool for Solving Equations @mcp.tool(name="solve_equation") def solve_equation(equation: str, variable: str): """ Solve equation for variable """ try: from sympy import sympify, symbols, solve, Basic def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x expr = sympify(equation) var = symbols(variable) res = solve(expr, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Solving Linear Systems @mcp.tool(name="solve_linear_system") def solve_linear_system(system: list, variables: list): """ Solve system of linear equations """ try: from sympy import sympify, symbols, linsolve, Basic def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x eqs = [sympify(e) for e in system] vars_sym = [symbols(v) for v in variables] res = linsolve(eqs, vars_sym) return {"success": True, "result": ser(list(res))} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Differentiation @mcp.tool(name="differentiate") def differentiate(expr: str, variable: str): """ Calculate derivative of expression with respect to variable """ try: from sympy import sympify, symbols, diff, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = diff(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Integration @mcp.tool(name="integrate_expression") def integrate_expression(expr: str, variable: str): """ Calculate integral of expression with respect to variable """ try: from sympy import sympify, symbols, integrate, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = integrate(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Polynomial Creation @mcp.tool(name="create_polynomial") def create_polynomial(expr: str, variable: str = None): """ Create polynomial from expression """ try: from sympy import sympify, symbols, Poly, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) if variable: var = symbols(variable) res = Poly(expr_sym, var) else: res = Poly(expr_sym) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Polynomial Factoring @mcp.tool(name="factor_polynomial") def factor_polynomial(poly: str): """ Factor polynomial expression """ try: from sympy import sympify, factor, Basic def ser(x): if isinstance(x, Basic): return str(x) return x poly_sym = sympify(poly) res = factor(poly_sym) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Fourier Transform @mcp.tool(name="fourier_transform") def fourier_transform(expression: str, time_var: str, freq_var: str): """ Calculate Fourier transform """ try: from calc.symbolic import sympify, symbols import numpy as np from scipy import integrate expr = sympify(expression) t = symbols(time_var) omega = symbols(freq_var) def integrand(t_val): return float(expr.subs(t, t_val)) t_range = (-50, 50) omega_vals = np.linspace(-10, 10, 100) result = [] for omega_val in omega_vals: def f(t_val): return integrand(t_val) * np.exp(-1j * omega_val * t_val) integral_result, _ = scipy_integrate.quad(f, t_range[0], t_range[1], limit=100) result.append(complex(integral_result)) return {"success": True, "result": [str(complex(r)) for r in result]} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Laplace Transform @mcp.tool(name="laplace_transform") def laplace_transform(expression: str, time_var: str, laplace_var: str): """ Calculate Laplace transform of a function. """ try: from calc.symbolic import sympify, symbols from calc.transforms import laplace_transform as sympy_laplace_transform import numpy as np from scipy import integrate expr = sympify(expression) t = symbols(time_var) s = symbols(laplace_var) try: F = sympy_laplace_transform(expr, t, s, noconds=True) return {"success": True, "result": str(F)} except Exception: def integrand(t_val): return float(expr.subs(t, t_val)) t_range = (0, 100) s_vals = np.linspace(0.1, 10, 50) result = [] for s_val in s_vals: def f(t_val): return integrand(t_val) * np.exp(-s_val * t_val) integral_result, _ = scipy_integrate.quad(f, t_range[0], t_range[1], limit=100) result.append({"s": float(s_val), "F": str(complex(integral_result))}) return {"success": True, "result": result} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Z Transform @mcp.tool(name="z_transform") def z_transform(expression: str, time_var: str, z_var: str, limit: int = 100): """ Calculate Z transform """ try: def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x expr = sympify(expression) n = symbols(time_var) z = symbols(z_var) try: result = Sum(expr * (z ** (-n)), (n, 0, $o$)).doit() result = simplify(result) return {"success": True, "result": ser(result)} except Exception: try: if expr.is_Pow and expr.base.is_symbol and expr.exp == n: a = expr.base result = z / (z - a) return {"success": True, "result": ser(result)} elif expr.is_Mul and len(expr.args) == 2: for arg in expr.args: if arg == n: other = expr / n if other.is_Pow and other.base.is_symbol and other.exp == n: a = other.base result = a * z / (z - a)**2 return {"success": True, "result": ser(result)} elif expr.is_Mul and n**2 in expr.args: other = expr / n**2 if other.is_Pow and other.base.is_symbol and other.exp == n: a = other.base result = a * z * (z + a) / (z - a)**3 return {"success": True, "result": ser(result)} result = 0 for k in range(limit + 1): term = expr.subs(n, k) * (z ** (-k)) result += term return {"success": True, "result": ser(result)} except Exception: result = 0 for k in range(limit + 1): term = expr.subs(n, k) * (z ** (-k)) result += term return {"success": True, "result": ser(result)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
A.3.3 Tools Generated from the Foam-Agent Repository
MCP Tool for Mesh Generation @mcp.tool(name="generate_mesh", description="Generate computational mesh using Foam-Agent internals.") def generate_mesh(requirements: str, case_dir: str = "./output", mesh_mode: str = "gmsh", custom_mesh_path: str | None = None): try: from src.config import Config from src.main import initialize_state from src.nodes.meshing_node import meshing_node config = Config() config.case_dir = case_dir state = initialize_state(user_requirement=requirements, config=config, custom_mesh_path=custom_mesh_path) if mesh_mode == "custom": state["mesh_type"] = "custom_mesh" elif mesh_mode == "gmsh": state["mesh_type"] = "gmsh_mesh" else: state["mesh_type"] = "standard_mesh" res = meshing_node(state) return {"success": True, "result": res, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Simulation Generation and Run @mcp.tool(name="generate_simulate", description="Write inputs and run simulation via Foam-Agent graph.") def generate_simulate(requirements: str, case_dir: str = "./output", custom_mesh_path: str | None = None, run_mode: str = "auto"): try: from src.config import Config from src.main import create_foam_agent_graph, initialize_state config = Config() config.case_dir = case_dir state = initialize_state(user_requirement=requirements, config=config, custom_mesh_path=custom_mesh_path) if custom_mesh_path: state["mesh_type"] = "custom_mesh" if run_mode == "local": state["cluster_info"] = None elif run_mode == "hpc": state["cluster_info"] = {"scheduler": "slurm"} workflow = create_foam_agent_graph().compile() workflow.invoke(state) return {"success": True, "result": {"case_dir": config.case_dir}, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Velocity Visualization @mcp.tool(name="visualize_velocity", description="Post-process and visualize velocity (|U|, streamlines, slices).") def visualize_velocity(case_dir: str, plot_type: str = "magnitude", plane: str | None = "xy"): try: from src.config import Config from src.main import initialize_state from src.nodes.visualization_node import visualization_node config = Config() config.case_dir = case_dir state = initialize_state(user_requirement="", config=config, custom_mesh_path=None) state["case_dir"] = case_dir state["visualization_request"] = {"plot_type": plot_type, "plane": plane} vis_res = visualization_node(state) return {"success": True, "result": vis_res, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}

### v3

@@
 1 Introduction
-The landscape of artificial intelligence is increasingly defined by autonomous agents that leverage Large Language Models (LLMs) to interact with external tools (Wang et al., 2024; Xi et al., 2024; Bubeck et al., 2023).
-To overcome the inherent limitations of LLMs in tasks requiring real-time information or precise computation, the paradigm of tool-augmented reasoning has become central (Huang et al., 2024; Hao et al., 2023; Yue et al., 2024).
+The landscape of artificial intelligence is increasingly defined by autonomous agents that leverage Large Language Models (LLMs) to interact with external tools (Wang et al., 2024; bubeck2023sparks).
+To overcome the inherent limitations of LLMs in tasks requiring real-time information or precise computation, the paradigm of tool-augmented reasoning has become central (Huang et al., 2024; Hao et al., 2023).
 Seminal works have demonstrated that models can effectively learn to invoke external functions (Schick et al., 2023; Qin et al., 2023; Parisi et al., 2022).
@@
 To address this, MCP is proposed as a universal standard that specifies how agents and tools should communicate, enabling an interoperable “plug-and-play” ecosystem (Anthropic, 2023).
-In response to this integration challenge, the community’s efforts have evolved, inadvertently revealing a deeper, more foundational bottleneck (Yue et al., 2025a).
-The initial challenge is to establish the fundamental feasibility of tool use, where a limited set of tools proves the feasibility of the paradigm (Schick et al., 2023; Patil et al., 2023; Ding et al., 2025).
-To break past the inherent scarcity of these platforms, the focus shifts to the vast landscape of open-source repositories (Wang et al., 2025; Xie et al., 2023).
+In response to this integration challenge, the community’s efforts have evolved, inadvertently revealing a deeper, more foundational bottleneck (Yue et al., 2025).
+The initial challenge is to establish the fundamental feasibility of tool use, where a limited set of tools proves the feasibility of the paradigm (Schick et al., 2023; Patil et al., 2023; Ding et al., 2025; Qin et al., 2024).
+To break past the inherent scarcity of these platforms, the focus shifts to the vast landscape of open-source repositories (Wang and others, 2025; Xie et al., 2023; Wang, 2023).
 This move, however, trades a scarcity problem for a chaos problem, exposing the wild non-standardization of real-world code.
-The MCP emerges as a direct answer to this chaos, promising a universal interface (Zhang et al., 2025; Anthropic, 2023).
-Yet, this leads to the critical gap: research is focused almost exclusively on the consumption side of MCP, using services from a presumed-to-exist pool (Gan & Sun, 2025), while the foundational “supply-side” problem of how to populate this pool from existing software is largely unaddressed.
+The MCP emerges as a direct answer to this chaos, promising a universal interface (Yu et al., 2025; Anthropic, 2023).
+Yet, this leads to the critical gap: research is focused almost exclusively on the consumption side of MCP, using services from a presumed-to-exist pool (Gan and Sun, 2025), while the foundational “supply-side” problem of how to populate this pool from existing software is largely unaddressed.
 Figure 1: While most research focuses on the consumption of tools (right side), one bottleneck is their supply (left side).
@@
 This supply bottleneck is not a theoretical concern but a stark reality preventing standards like MCP from achieving widespread adoption.
-For example, RAG-MCP (Gan & Sun, 2025) utilizes over 4,400 servers on mcp.so, but there are 268 million public GitHub repositories.
+For example, RAG-MCP (Gan and Sun, 2025) utilizes over 4,400 servers on mcp.so, but there are 268 million public GitHub repositories.
 The critical question of how to create a large and diverse pool of these standardized, agent-ready tools has been left unaddressed.
@@
 To systematically address these challenges, as shown in Figure 2, Code2MCP implements a collaborative multi-agent system (Park et al., 2023).
-Unlike general-purpose coding agents (Cognition, 2024; Jimenez et al., 2024), different agents in our framework are specialized for the distinct stages of code analysis, environment setup, and API design.
+Unlike general-purpose coding agents (Cognition, 2024; Jimenez et al., 2024; Zhang et al., 2024b; Huang et al., 2023), different agents in our framework are specialized for the distinct stages of code analysis, environment setup, and API design.
 Crucially, the overall reliability of this workflow is ensured by an integrated Run-Review-Fix self-correction cycle, which endows the system with the ability to autonomously debug and refine the entire conversion process.
 The key contributions of this work are listed as follows:
-  • To solve the fundamental tool supply bottleneck hindering the adoption of the MCP standard, we propose a novel automated framework, Code2MCP, the first framework to systematically transform code repositories into agent-ready MCP services.
+  • To solve the fundamental tool supply bottleneck hindering the adoption of the MCP standard, we propose Code2MCP, a novel automated framework that, to the best of our knowledge, is the first to systematically transform code repositories into agent-ready MCP services.
   • The key challenge in converting code into a service is the inherent fragility of the multi-stage automation process, where an error at any step can derail the entire workflow.
@@
 Work Core Contribution Tool Pool Tool Selection MCP
-Consumer
-Toolformer Teaching LLM to use external tools 5 predefined tools Fine-tuning $\times$
-SciToolAgent
-Domain-specific enhancement
-for scientific tool utilization
-KG of Scientific Tools
-(500+ tools)
-Retrieval on KG $\times$
-HuggingGPT
-Increase the size
-of tool pool
-huggingface.co LLM task planning $\times$
-Gorilla
-TorchHub, TensorHub
-(1600+ tools)
-Retriever-aware
-training
-$\times$
-OpenAgents
-Tool use from open-source
-beyond a closed pool
-github.com Multi-agent planning $\times$
+Consumer Toolformer Teaching LLM to use external tools 5 predefined tools Fine-tuning $\times$
+SciToolAgent Domain-specific enhancement for scientific tool utilization KG of Scientific Tools (500+ tools) Retrieval on KG $\times$
+HuggingGPT Increase the size of tool pool huggingface.co LLM task planning $\times$
+Gorilla TorchHub, TensorHub (1600+ tools) Retriever-aware training $\times$
+OpenAgents Tool use from open-source beyond a closed pool github.com Multi-agent planning $\times$
 RepoMaster Rule-based deep search $\times$
@@
 A more ambitious paradigm shift involves treating the entirety of open-source code repositories as a virtually infinite tool source.
-Frameworks like OpenAgents (Xie et al., 2023) and RepoMaster (Wang et al., 2025) empower agents to directly parse, reason about, and execute code within GitHub repositories.
+Frameworks like OpenAgents (Xie et al., 2023) and RepoMaster (Wang and others, 2025) empower agents to directly parse, reason about, and execute code within GitHub repositories.
 These works confront the complexity of real-world code but expose the core bottleneck: the vast majority of these repositories are not designed for programmatic use by LLM agents.
-They lack standardized interfaces (Zhang et al., 2024; Jin et al., 2024; Ray, 2025), forcing the agent into an ad-hoc, brittle, and unreliable process of reverse-engineering the code, setting up its environment, and managing dependencies for every single task (Zeng et al., 2024; Olausson et al., 2023).
+They lack standardized interfaces (Zhang et al., 2024a; Jin et al., 2024; Ray, 2025), forcing the agent into an ad-hoc, brittle, and unreliable process of reverse-engineering the code, setting up its environment, and managing dependencies for every single task.
 This chaotic integration process illustrates a critical failure on the tool supply side.
@@
 Recognizing this chaos, the community has moved towards standardization, exemplified by the MCP.
-For example, RAG-MCP (Gan & Sun, 2025) explores how an agent can effectively retrieve and select the most appropriate MCP service from a pool of 4400+ available options.
+For example, RAG-MCP (Gan and Sun, 2025) explores how an agent can effectively retrieve and select the most appropriate MCP service from a pool of 4400+ available options.
 This approach is promising, but it presumes the existence of a rich ecosystem of MCP-compliant services (Hasan et al., 2025).
@@
 The entire conversion process, as depicted in Figure 2, is a multi-agent workflow that begins with code analysis, proceeds through a core Run-Review-Fix self-correction loop, and culminates in the generation of a merge-ready pull request.
+For completeness, a more formal description of this workflow is provided in the appendix A.3.
 Suppose there exists a consumer-side work listed in Table 1 that finds a suitable GitHub repository that may solve the user’s query.
@@
 This is the core difference between this “supply-side” work and the consumer-side works.
+Initialization and Analysis.
+The Download Agent first clones the specified repository, identified by its URL $u$, into an isolated local workspace.
+The Environment Agent then replicates the runtime environment from dependency files or Dockerfiles, addressing one of the most common failure points in code conversion and supporting reliable subsequent code generation and testing.
+Once the environment is ready, the Analysis Agent identifies tool-worthy functionalities within the codebase.
+It leverages the DeepWiki tool to obtain a semantic view of the code and associates code entities with their intent from documentation and comments.
+The output is a Code Report that summarizes candidate APIs and guides subsequent stages.
+Generation, Execution, and Self-Correction. Given this conversion blueprint, the framework enters its core iterative loop that turns the identified functionalities into executable MCP services.
+The loop starts with the Generation Agent, which takes the Code Report and uses an LLM to abstract the core functionalities into MCP-compliant interfaces.
+It creates the tool interface definitions and adapter file that connect the original code to the MCP interface, together with a basic test suite.
+Once the code is generated, the Run Agent executes the test suite in the prepared environment to verify executability.
+If the tests pass, the workflow proceeds to finalization; otherwise, the Run Agent records the error traceback $\tau$ and forwards it to the Review Agent.
+The Review Agent analyzes $\tau$ together with the generated code, the Code Report, and the failing test, and diagnoses root causes such as logic errors, missing dependencies, or interface mismatches.
+It then formulates a correction plan $\delta$ that specifies which files and code blocks to change, and hands this plan back to the Generation Agent to re-synthesize the MCP files.
+This Run-Review-Fix loop repeats until tests pass or a maximum of $B$ attempts is reached.
+Finalization and Delivery.
+After the core loop succeeds, the Finalize Agent organizes and packages the validated MCP service files.
+To facilitate review and adoption by the original repository maintainers, it also generates a README file explaining how to use the new MCP service.
+All artifacts are arranged into a reproducible directory structure under the workspace (Appendix A.2), and the agent can prepare a pull request to submit these additions back to the original repository.
+4 Evaluation
+4.1 Experimental Setup
+Task.
+The evaluation task is: given a GitHub repository URL, Code2MCP automatically generates an MCP service for that repository, and the outcome of the conversion is judged as a success or a failure.
+The evaluation focuses on the overall conversion success rate across different domains and repository characteristics, and on the time and stability of Code2MCP compared with manual implementation and a GPT-4 template-based baseline on the same set of repositories.
+Repositories.
+Code2MCP is evaluated on 50 open-source repositories from 10 scientific and engineering domains (5 per domain), covering biomedical science, psychology, mathematical computing, earth science, chemistry, physics, astronomy, social science, linguistics, and econometrics.
+These repositories range from small utility libraries to large frameworks with complex dependency stacks.
+For each repository, the evaluation logs the environment construction result, MCP conversion result, number of generated tools, and failure types observed across all Run-Review-Fix rounds; per-repository details and the associated failure labels are summarized in a large table in the Appendix C, and a subset of representative scientific-computing repositories is further used for the Run-Review-Fix ablation and qualitative case studies.
+Success criterion.
+A conversion is regarded as successful only if the following three conditions all hold: the runtime environment can be reconstructed according to the dependency specifications provided by the repository, the generated MCP server can be started and must pass a unified RunNode test, and at least three core tools can be invoked correctly in automatic tests and return valid outputs.
+If any of these conditions is not satisfied, the repository is counted as a failure.
+Failure taxonomy. To analyze failure modes, all failed repositories are annotated with one or more labels from the following six categories:
+  • env_failure. The environment or dependencies cannot be reconstructed.
+  • api_inference_error. Systematic errors occur when inferring tool interfaces.
+  • import_error. Import paths or cross-module dependencies are misconfigured.
+  • repo_internal_bug. Bugs or version conflicts inside the original repository prevent it from running reliably even in its native environment.
+  • mcp_spec_violation. The generated JSON Schema or response structure does not conform to the MCP specification.
+  • untoolable_repo. The repository is a collection of scripts or heavily depends on interactive CLIs or GUIs, which are difficult to abstract into stable MCP tools.
+Implementation Details.
+By default, Code2MCP utilizes gpt-4o-2024-05-13 as its core reasoning engine.
+The temperature for all models is set to 1.
+Code2MCP leverages gitingest [^1] https://github.com/coderamp-labs/gitingest to ingest repositories into contextual prompts and fetch pre-analysis reports from deepwiki [^2] https://deepwiki.org.
+Case studies are conducted on servers equipped with 8 NVIDIA H100 80 GB GPUs.
+Table 2: Per-domain summary of environment setup success, basic test success, recovery by the Run-Review-Fix (RRF) loop (“–” if none are recovered), and final MCP conversion success.
+Domain Repos Env succ. Test succ. RRF recovered Avg rounds MCP succ.
+Biomedical 5 4/5 (80%) 2/5 (40%) 1/2 (50%) 1.5 3/5 (60%)
+Psychology 5 3/5 (60%) 3/5 (60%) 0/1 (0%) 3/5 (60%)
+Math 5 5/5 (100%) 3/5 (60%) 1/2 (50%) 2.0 4/5 (80%)
+Earth Science 5 3/5 (60%) 2/5 (40%) 1/3 (33.3%) 1.0 3/5 (60%)
+Chemistry 5 4/5 (80%) 2/5 (40%) 1/2 (50%) 1.3 3/5 (60%)
+Physics 5 3/5 (60%) 2/5 (40%) 0/2 (0%) 2/5 (40%)
+Astronomy 5 4/5 (80%) 3/5 (60%) 1/1 (100%) 1.5 4/5 (80%)
+Social Science 5 4/5 (80%) 2/5 (40%) 1/2 (50%) 1.0 3/5 (60%)
+Linguistics 5 5/5 (100%) 3/5 (60%) 1/2 (50%) 2.0 4/5 (80%)
+Econometrics 5 3/5 (60%) 2/5 (40%) 1/2 (50%) 1.0 3/5 (60%)
+Overall 50 38/50 (76%) 24/50 (48%) 8/19 (42.1%) 1.4 32/50 (64%)
+4.2 Large-scale Repository Conversion and Failure Analysis
+Overall and per-domain success.
+The overall conversion success rate is first analyzed at the domain level.
+In Table 2, for each domain, the number of repositories with successful environment setup and basic test passing, the number of initially failing repositories recovered by the Run-Review-Fix loop and their average rounds, and the final number of successful MCP conversions.
+The per-domain MCP success rates range from 0.40 to 0.80: domains dominated by library-style projects, such as mathematical computing, astronomy, and linguistics, tend to have higher success rates, while domains such as physics and econometrics, which contain more complex environments and mixtures of scripts and workflow-style code, tend to have lower success rates.
+All statistics in the table are computed on a per-repository basis, where each repository contributes a single final outcome after the Run-Review-Fix loop.
+Per-repository results, including environment construction and MCP conversion outcomes as well as the number of generated tools, are summarized in the large table in the appendix and can be further grouped by domain or repository size if desired.
+Figure 3: Distribution of 42 failure labels assigned to the 18 failed repositories. A single repository can trigger multiple failure types across different Run-Review-Fix rounds.
+Failure modes.
+To understand why conversions fail on real repositories, all 18 failed repositories in the 50-repository evaluation are annotated using the six failure labels introduced in the failure taxonomy.
+Labels are aggregated across all Run-Review-Fix rounds, so a single repository may receive multiple labels if different error types are observed during different stages of the pipeline.
+The distribution of these labels is visualized in Figure 3.
+The labels env_failure and api_inference_error account for more than half of all failures, indicating that environment reconstruction and interface inference are the primary bottlenecks of the current workflow.
+The labels import_error and untoolable_repo are also relatively common, reflecting the difficulty of handling complex import paths, initialization order, and script-driven repositories in a fully automated pipeline.
+Run-Review-Fix ablation.
+To assess the specific contribution of the Run-Review-Fix self-correction loop, an ablation study is conducted on a set of representative scientific-computing repositories by comparing two configurations: a single-pass configuration without the loop and the full Code2MCP pipeline with the loop enabled.
+Figure 4 summarizes the results.
+The bar chart on the left shows that the per-repository success rates increase for most repositories once the loop is enabled, often moving from medium success under a single-generation setting to success rates close to or on par with human-written wrappers.
+The line chart on the right plots the average number of remaining errors as a function of the number of Run-Review-Fix rounds, where “remaining errors” denote the number of failing assertions or uncaught exceptions in the unified test after each round.
+These results indicate that the majority of initial failures can be automatically repaired within one to three Run-Review-Fix rounds, whereas the few repositories that still fail after multiple rounds represent the main current limitation of the method.
+4.3 Comparison with Human and GPT-4 Baselines
+The evaluation compares three configurations on the same set of representative repositories:
+  • a human configuration, where MCP wrappers are implemented from scratch by human developers;
+  • a GPT-4 template configuration, where a single GPT-4 agent generates MCP wrappers;
+  • the Code2MCP configuration, which uses the multi-agent pipeline with the Run-Review-Fix loop.
+Figure 4: Repo success rates with and without Run‑Review‑Fix (left) and remaining errors across RRF rounds (right) for five representative scientific MCP repositories.
+In the human configuration, ten graduate students with at least three years of Python experience and basic familiarity with MCP concepts implement wrappers from scratch.
+Before the study, all participants read the MCP specification and a commented example repository.
+Each participant is randomly assigned several repositories, and the measured time is from the moment they start reading the repository to the point where the minimal test passes.
+In the GPT-4 template configuration, a unified system prompt instructs a single GPT-4 model to generate a complete MCP service implementation based on the repository README and several key source files selected by static analysis.
+We use the same base model and decoding configuration as in Code2MCP, namely gpt-4o-2024-05-13 with temperature set to 1.0, to ensure a fair comparison.
+The model is allowed up to three dialogue turns: the first turn generates an initial implementation, and up to two additional turns may revise the code based on execution errors returned by a test runner.
+No additional agents or planning mechanisms are used, and the full system prompt is provided in Appendix B.1.
+Figure 5: Average task success rate (left) and average completion time (right) across the ten scientific domains for the three configurations: Human experts, GPT-4 with basic scaffolding, and Code2MCP.
+In the Code2MCP configuration, the default multi-agent pipeline with the Run-Review-Fix loop is used.
+For each repository, we record the final success status and the completion time, defined as the end-to-end wall-clock time starting from the beginning of environment setup and ending when the unified MCP test first passes or the run is terminated; see Appendix B.2 for details.
+Figure 5 summarizes the average task success rates and completion times of the three configurations across the ten scientific domains.
+For each configuration and domain, task success rate is defined as the fraction of successful attempts over all attempts on representative repositories in that domain.
+Completion time is the average end-to-end wall-clock time per attempt, measured from the beginning of environment setup until the minimal MCP test first passes or the run is terminated.
+Figure 6: Task success rates of Human experts, GPT-4 with a template, and Code2MCP on representative repositories in the four repository groups.
+The x-axis lists representative repositories in each group, labeled by domain abbreviation and index (e.g., Math-1).
+The overall trend is as follows: Code2MCP achieves task success rates close to human implementations in most domains, and consistently higher than the GPT-4 template configuration.
+In terms of completion time, human implementations are typically on the order of hours, while Code2MCP completes in tens of minutes; the GPT-4 template configuration is the fastest but has lower success rates.
+When repositories are grouped by dependency complexity and project type into low-dependency libraries, medium-dependency scientific libraries, high-dependency libraries, and script-heavy projects, Figures 6 and 7 further break down the same comparison on representative repositories within each group.
+The bucketing rules are detailed in Appendix B.2.
+Figure 7: Completion time of Human experts, GPT-4 with a template, and Code2MCP on representative repositories in the four repository groups.
+4.4 Integration with Existing Tool Systems
+This subsection summarizes how Code2MCP integrates with existing tool ecosystems by combining it with RepoMaster and OpenAgents on a shared set of tool-usage tasks.
+RepoMaster + Code2MCP.
+In the RepoMaster setting, code-understanding and tool-usage tasks are run in two configurations: using RepoMaster alone to operate at the source-code level, and using RepoMaster together with Code2MCP, where RepoMaster first selects relevant repositories and then invokes the corresponding MCP tools.
+Figure 8 shows that the configuration with Code2MCP achieves higher task success rates and requires fewer interaction steps (see Appendix B.2 for metric definitions), suggesting that delegating part of the work to standardized MCP tools improves efficiency without modifying the upstream agent policy.
+OpenAgents + Code2MCP.
+In the OpenAgents setting, cross-domain tasks are evaluated under two configurations: using OpenAgents alone, and using OpenAgents with MCP tools automatically generated by Code2MCP added to the tool pool.
+As shown in Figure 9, we report the task success rate and coverage, where coverage and success rate are defined in Appendix B.2.
+Adding Code2MCP tools increases both metrics across all task groups, indicating that a richer MCP tool pool can be effectively exploited by existing planning and retrieval strategies.
+Overall, RepoMaster and OpenAgents handle tool discovery and code retrieval on the consumer side, while Code2MCP supplies additional MCP services on the supply side.
+Under the same front-end selection and planning strategies, integrating Code2MCP enriches the tool pool and yields measurable improvements in end-to-end task performance.
+Figure 8: RepoMaster with and without Code2MCP tools: task success rate (left) and average interaction steps (right) across task groups.
+4.5 Case Studies: Protein, Math, and Computational Fluid Dynamics
+This subsection highlights three representative repositories from the 50-repository evaluation set: the biomedical protein modeling library ESM, the symbolic mathematics library SymPy, and the CFD framework Foam-Agent built on OpenFOAM, illustrating the kinds of MCP tools Code2MCP generates and how agents use them in practice.
+Figure 9: OpenAgents with and without Code2MCP tools: task success rate (left) and coverage (right) across general and scientific task groups.
+Bridging protein science with AI agents.
+In protein science, models such as AlphaFold (Jumper et al., 2021) greatly improve structure prediction, but using them in everyday research typically requires substantial scripting and environment setup.
+ESM models complement AlphaFold with efficient sequence modeling and zero-shot variant effect prediction, yet traditionally still require multi-step Python pipelines for computing physicochemical properties, predicting structures, and analyzing mutations.
+Code2MCP converts ESM into an MCP service by exposing tools such as Analyze_sequence, Predict_structure, and Predict_variant_effect.
+We provide a qualitative example of using ESM-based MCP tools for protein analysis in Appendix A.5.
+Figure 10: An AI agent orchestrates a CFD simulation pipeline by invoking Generate_mesh, Generate_simulate, and Visualize_velocity functions to guide a user through the entire process, from initial mesh generation and simulation setup to the final velocity visualization.
+Enhancing mathematical reliability in AI agents.
+For mathematical computing, large language models are prone to errors in symbolic derivations and exact calculations, whereas libraries such as SymPy already provide reliable implementations.
+Code2MCP organizes common capabilities from SymPy into a mathematical MCP service with tools for limits and integration, matrix operations, symbolic simplification, and transforms such as Fourier and Laplace.
+When asked to compute an integral or the volume of a solid of revolution, an agent simply invokes the corresponding tool and returns an exact answer, relying on SymPy as a backend instead of ad-hoc prompt-based reasoning.
+Automating CFD simulation for AI agents.
+CFD workflows typically involve mesh generation, solver configuration, execution, and post-processing, traditionally coordinated by engineers through shell scripts and configuration files.
+Foam-Agent provides a higher-level interface to OpenFOAM, and Code2MCP further turns it into an MCP service with tools such as Generate_mesh, Generate_simulate, and Visualize_velocity.
+As illustrated in Figure 10, a user describes the target geometry and physical parameters in natural language, and the agent calls these tools in sequence to construct the mesh, run the simulation, and visualize the velocity field.
+5 Conclusion
+Addressing the key challenge of insufficient tool supply in the MCP ecosystem, this paper introduces Code2MCP, a framework that automatically converts GitHub repositories into functional MCP services.
+The framework employs a multi-agent workflow for code comprehension, environment reconstruction, and tool abstraction, augmented by a “Run-Review-Fix” self-correction loop that improves end-to-end reliability across diverse scientific libraries.
+Our evaluation on 50 repositories shows that Code2MCP can robustly expose high-value APIs as MCP tools while also revealing clear failure modes and boundary conditions.
+This work presents an automated pathway for connecting existing codebases to the agent tool ecosystem, and we leave systematically hardening security and building standardized large-scale MCP benchmarks from these converted services as important directions for future work.
+Appendix A Appendix
+A.1 Use of Large Language Models
+During manuscript preparation, large language models (LLMs) are used solely as general-purpose writing assistants for grammar checking, wording refinement, and improving clarity.
+LLMs don’t contribute to research ideation, methodological design, or experimental execution.
+All suggestions produced by LLMs are reviewed, edited, and vetted by the authors, who take full responsibility for the entire content of the paper.
+A.2 Output Directory Structure
+Figure 11 shows the final directory layout generated by Code2MCP, where the original repository and the synthesized MCP artifacts are organized under a unified workspace.
+Figure 11: The complete output directory structure generated by the Code2MCP framework.
+The top-level workspace contains the original repository alongside all generated artifacts within the mcp_output directory.
+A.3 Detailed Conversion Pipeline and Algorithm
+This section outlines the roles of the specialized agents within the Code2MCP framework and their coordination as shown in Algorithm 1, where Download and Environment initialize the workspace, Analysis and Generation construct the MCP service, Run and Review form the Run-Review-Fix loop, and Finalize produces the final deliverables.
+The following are placeholders for their system prompts.
 Algorithm 1 The Code2MCP Framework
@@
 11:  else
-12:   Review Agent: Analyze $\tau$ $\rightarrow$ generate correction plan $\delta$.
-13:   Generation Agent: Re-synthesize MCP files using Code Report and correction plan $\delta$.
+12:   Review Agent: Analyze traceback $\tau$ and generate correction plan $\delta$.
+13:   Generation Agent: Re-synthesize MCP files using the Code Report and correction plan $\delta$.
 14:   $r\leftarrow r+1$
@@
 16:end while
-17:Finalize Agent: Package service files, generate README, and create Pull Request.
+17:Finalize Agent: Package service files, generate README, and create a Pull Request.
 18:Output: A merge-ready Pull Request containing the functional MCP service.
-Initialization and Analysis.
-First, the Download Agent is responsible for cloning the specified repository, identified by its URL $u$, and creating an isolated local workspace.
-Subsequently, the Environment Agent begins its work by addressing environmental dependencies, one of the most common failure points in code conversion.
-It meticulously replicates the original code’s runtime environment by intelligently parsing dependency files or Dockerfiles.
-This step is the cornerstone for ensuring the reliability of all subsequent code generation and testing.
-Once the environment is ready, the core challenge is identifying valuable, tool-worthy functionalities within a complex codebase.
-This is the task of the Analysis Agent.
-It moves beyond traditional static analysis by leveraging the DeepWiki tool for a deep semantic understanding of the code.
-By associating code entities with their intent from documentation and comments, the agent understands the code’s “purpose”, not just its “structure”.
-The final output is a detailed Code Report, which serves as the strategic blueprint for all subsequent stages.
-Generation, Execution, and Self-Correction.
-After obtaining a clear conversion blueprint, the framework enters its core iterative loop, which is designed to tackle the challenge of translating abstract functionalities into concrete, executable, and correct MCP services.
-The loop commences with the Generation Agent.
-It takes the Code Report from the Analysis Agent and uses the code generation capabilities of an LLM to intelligently abstract and encapsulate the identified core functionalities into MCP-compliant interfaces.
-This includes creating essential files like the tool interface definition and adapter file, which connect the original code to the MCP interface, along with a basic test suite.
-Once the code is generated, the Run Agent immediately executes the test suite within the prepared environment.
-Its role is to verify the practical executability of the generated code.
-If the tests pass, the workflow exits the loop and proceeds to the finalization stage.
-However, if execution fails, the Run Agent captures a detailed error traceback, denoted as $\tau$, and passes it to the Review Agent, triggering the framework’s critical self-correction mechanism.
-The Review Agent acts as a “senior debugging engineer”, performing an in-depth error analysis.
-Its analysis goes beyond simply parsing the traceback $\tau$; it correlates the error message with the generated code, the original repository’s structure outlined in the Code Report, and the specific test case that failed.
-This contextual reasoning allows it to diagnose the root cause with high precision, identifying issues such as code logic error, a missing dependency, or an interface mismatch.
-After diagnosis, it formulates a correction plan, $\delta$.
-The resulting plan is a structured, actionable directive that specifies the exact file and code block to be modified and provides a clear description of the required change.
-This precise instruction is then fed back to the Generation Agent to re-synthesize the corrected code.
-This closed loop of “run-review-fix,” formed by the Run Agent and Review Agent, iterates until all issues are resolved and the generated code passes all tests.
-To prevent infinite loops, the process is bounded by a maximum of $B$ attempts.
-This robust self-correction mechanism is the key to Code2MCP’s end-to-end reliability.
-Figure 3: The complete output directory structure generated by the Code2MCP framework.
-The top-level workspace contains the original repository alongside all generated artifacts within the mcp_output directory.
-Finalization and Delivery.
-Upon the successful completion of the core loop, the Finalize Agent handles the concluding tasks.
-It organizes and packages all the validated, functional MCP service files.
-To facilitate review and adoption by the original repository maintainers, it also automatically generates a clear README file explaining how to use the newly added MCP service.
-Finally, the Finalize Agent packages all generated artifacts into the clear directory structure shown in Figure 3.
-Operating as a fully automated contributor, it then uses the GitHub API to submit these additions as a new Pull Request to the original repository.
-This act completes the code-to-service transformation and builds a direct bridge connecting the open-source project to the growing MCP ecosystem.
-4 Evaluation
-4.1 Experimental Setup
-Task.
-The evaluation task is to convert diverse GitHub repositories into validated MCP services, assessing both the efficiency in reducing conversion time and cost compared to manual effort, and the effectiveness demonstrated by the functional success of the generated tools.
-Demonstrated Tools.
-As shown in Table 2, we demonstrate 6 representative open-source repositories to cover a spectrum of complexity and domains.
-The first group includes three widely-used utility libraries for simpler conversion tasks: TextBlob for natural language processing, dateutil for datetime parsing, and searchkick for intelligent search.
-These serve to demonstrate the significant speedup Code2MCP provides for standard conversion tasks.
-The second group includes three complex scientific libraries to test Code2MCP’s ability to handle domain-specific codebases: ESM for protein science, SymPy for symbolic mathematics, and Foam-Agent for computational fluid dynamics.
-Their primary purpose is to serve as the subjects of our in-depth case studies.
-These cases provide the qualitative evidence for its success on complex software.
-This capability is crucial for addressing the tool supply bottleneck.
-Evaluation.
-We evaluate the effectiveness and generality of Code2MCP through a two-pronged approach: quantitative evaluation and in-depth case studies.
-The quantitative evaluation measures key performance metrics, such as conversion time and LLM token cost, against the estimated time for manual conversion to validate Code2MCP’s efficiency.
-Implementation Details.
-By default, Code2MCP utilizes gpt-4o-2024-05-13 as its core reasoning engine.
-The temperature for all models is set to 1.
-Code2MCP leverages gitingest [^1] https://github.com/coderamp-labs/gitingest to ingest repositories into contextual prompts and fetch pre-analysis reports from deepwiki [^2] https://deepwiki.org.
-Case studies are conducted on servers equipped with 8 NVIDIA H100 80 GB GPUs.
-Table 2: Statistics on the time required for Code2MCP to convert code into MCP and the token consumption. The manual time is an average of estimates from three senior developers.
-Repository Manual Time Code2MCP Speedup LLM Cost
-(Est. hours) Time (min) Factor (Tokens)
-sloria/TextBlob 1.0 – 1.5 12 $\sim$6x 24.7k
-dateutil/dateutil 0.7 – 1.2 15 $\sim$4x 19.2k
-ankane/searchkick 1.0 – 1.5 13 $\sim$6x 28.9k
-facebookresearch/esm 4.0 – 5.0 50 $\sim$5x 58.5k
-sympy/sympy 2.5 – 3.5 32 $\sim$6x 78.3k
-csml-rpi/Foam-Agent 2.0 – 3.0 17 $\sim$8x 41.4k
-4.2 Quantitative Evaluation
-A core motivation of Code2MCP is to alleviate the significant manual effort of turning code repositories into MCP services.
-Therefore, we report the time and LLM tokens consumed by Code2MCP in Table 2 and compare this with the estimated manual effort for a human expert.
-The manual time is estimated by three senior developers familiar with the respective libraries and the MCP standard, and the average value is reported.
-The manual task includes code comprehension, code wrapping, protocol implementation, testing, and documentation.
-These tasks are tedious and time-consuming, posing a significant challenge even for experienced developers.
-From Table 2, we observe that the estimated manual time far exceeds that of Code2MCP.
-This disparity arises because the manual task is a complex endeavor, encompassing challenges like deep code comprehension and environment replication, which extend far beyond simple implementation and testing.
-Having established the framework’s general efficiency, we now proceed to the case studies of the three scientific libraries.
-These studies on ESM, SymPy, and Foam-Agent will provide a qualitative analysis of the framework’s core capabilities.
-4.3 Case Studies: Protein, Math, and Computational Fluid Dynamics
-Figure 4: An AI agent processes a user’s query containing a protein sequence and mutation.
-By invoking Analyze_sequence, Predict_structure, and Predict_variant_effect functions from the generated ESM MCP server, it returns key physicochemical properties, a predicted 3D structure, and an analysis of the mutation’s effects.
-Bridging Protein Science with AI Agents.
-Advanced computational models, exemplified by AlphaFold (Jumper et al., 2021), have revolutionized structural biology with breakthroughs in protein structure prediction and sequence analysis.
-However, their practical application is hindered by steep learning curves and substantial engineering effort, creating a high barrier to entry for many scientists, especially those without deep computational backgrounds.
-For molecular biologists or drug discovery researchers focused on wet-lab experiments, writing Python scripts to analyze protein variants is often a significant distraction from their research.
-While AlphaFold is now present in the MCP ecosystem  [^3] https://github.com/augmented-nature/alphafold-mcp-server, ESM provides complementary capabilities such as zero-shot variant effect prediction to assess how mutations impact proteins, and it offers higher efficiency and accuracy.
-Thus, we select ESM as our conversion target.
-More broadly, scientific discovery depends on a diverse toolbox rather than a single universal solution, and the current scarcity of tool variety creates a critical “supply-side” bottleneck.
-Historically, performing even foundational protein characterization with a library like ESM is a multi-step, technical endeavor.
-Users have to write custom scripts, navigate the library’s internal structure to call the correct functions, and manually parse the outputs.
-Visualizing the data often necessitates integrating separate third-party packages as well.
-This entire process is demanding of programming expertise, tedious, and error-prone, ultimately yielding key physicochemical properties.
-In contrast, Code2MCP fully automates this entire process.
-The conversion itself is efficient, completed in about fifty minutes on a standard laptop.
-The resulting service hides the underlying technical complexity, making the tool accessible even to teams without programming expertise.
-Code2MCP encapsulates the repository’s core capabilities into a powerful ESM MCP tool containing ten composable functions.
-These functions support a range of tasks, from foundational analysis to interpreting the structural effects of variants.
-As shown in Figure 4, when a user provides a protein sequence and a mutation site like “Q145G”, the AI agent invokes several functions from the tool.
-First, it calls the Analyze_sequence function to provide a baseline characterization, returning the protein’s key physicochemical properties such as molecular weight, amino acid composition, and hydrophobicity.
-Next, it calls the Predict_structure function to generate the protein’s 3D structure and uses a Predicted Aligned Error (PAE) plot to assess its reliability.
-Finally, the agent uses the Predict_variant_effect function.
-This function not only identifies the mutation’s location on the 3D structure and analyzes its potential impacts, such as disrupting local stability or altering ligand interactions, but it can also generate a 2D projection map to visually clarify the mutation’s spatial context and its proximity to key functional regions.
-By integrating the outputs from these functions, the agent provides a complete analysis, clearly outlining the entire process from sequence variation to structural change and, finally, to functional impact.
-This comprehensive analysis directly connects a protein variation to its potential structural and functional consequences, a critical step in fields like personalized medicine or protein engineering.
-Code2MCP enables an AI agent to manage this process end-to-end, delivering immediate and interpretable results without requiring code from the user.
-This automated conversion and simplified interaction reduces setup and runtime, allowing researchers to focus on scientific questions rather than code logistics.
-Enhancing Mathematical Reliability in AI Agents.
-We next present a case study of Code2MCP in mathematics, where we create a comprehensive math MCP service by converting several core libraries, including SymPy.
-This case reflects a common challenge in science and engineering: LLMs are unreliable for precise computation.
-Their answers are difficult to verify, and minor errors can lead to incorrect conclusions.
-This forces agents to rely on their own numerical processing, making complex calculations a significant source of error and hallucination and undermining user trust in scientific and engineering applications.
-Meanwhile, powerful and validated codebases like SymPy are isolated from the modern AI ecosystem.
-Leveraging these libraries historically requires users to write custom Python scripts and manually translate mathematical problems using library’s specific functions and structures, a process that is both technically demanding and inefficient.
-Using just the links to these math GitHub repositories, Code2MCP automatically generates a fully featured MCP service.
-This new service comprises 25 atomic functions covering a wide range of capabilities from symbolic algebra, calculus, and matrix operations to probability, statistics, and various mathematical transforms.
-Figure 5 shows a selection of these tools.
-For a specific problem, such as finding the indefinite integral of $y=\ln(x^{2}+1)$, an agent can invoke the newly added Integrate_expression function.
-Similarly, for a task like finding the volume of revolution for $y=\ln(x)\cdot\sin(x)$ from $x=1$ to $x=2\pi$, it uses the converted Calculate_volume function, producing an output that matches the result from an expert manually using the original library.
-For more advanced research, an agent can leverage this suite of new functions for specialized, domain-specific calculations, extending its capabilities beyond simple arithmetic.
-This elevates the agent from a general-purpose assistant to a domain-aware collaborator capable of executing complex scientific workflows with verifiable precision.
-For example, a physicist can use the Laplace_transform function to analyze the stability of control systems or electrical circuits, while a digital signal processing expert can use the Z_transform function to analyze discrete-time signals for applications like designing audio filters.
-Furthermore, a financial analyst can employ the Fourier_transform to identify cyclical patterns in market data, thus unlocking new, specialized domains for AI agents by providing the validated functions they need.
-Figure 5: The figure shows examples from the service’s mathematical toolkit, organized into three functional categories: “Value” for numerical calculus operations like calculating limits and areas; “Symbol” for symbolic algebra such as simplifying and integrating expressions; “Transform” for advanced operations like Fourier and Laplace transforms.
-Case Study: Automating CFD Simulation for AI Agents.
-While the previous case studies focus on encapsulating individual functions, this section tackles the more profound challenge of automating multi-stage engineering workflows.
-Computational Fluid Dynamics (CFD) serves as a prime example of such a workflow.
-Traditionally, its interdependent steps require meticulous manual intervention.
-An engineer must write detailed scripts for tasks like geometry definition and manually edit complex text-based files to set solver parameters.
-This process is not only tedious and prone to syntax errors but also demands deep expertise in the repository’s internal structure.
-We selected Foam-Agent, a repository specifically for CFD simulations (Yue et al., 2025b).
-It is a framework designed to streamline operations within the OpenFOAM library (Jasak, 2009), capable of handling traditional tasks in CFD simulations such as mesh generation, solver configuration, and post-processing.
-However, its inherent execution logic follows a rigid, end-to-end sequential process, which lacks the flexibility required for dynamic adjustments or interrupt-resume scenarios in practical applications.
-Code2MCP overcomes this rigidity by automatically decomposing Foam-Agent’s integrated simulation process into a set of independent atomic functions, such as Generate_mesh, Generate_simulate, and Visualize_velocity.
-This allows the agent to intelligently coordinate the entire complex simulation procedure through multi-turn dialogue with a user (Schick et al., 2024).
-This approach replaces tedious manual configuration file editing and transforms a command-line program into a flexible, interactive scientific partner.
-For example, the entire CFD simulation workflow, beginning with a general instruction such as “simulate flow over a cylinder” followed by the configuration of specific physical parameters, and concluding with the final request to “visualize the velocity field”, can be completed step-by-step through natural language dialogue between the user and the agent.
-As illustrated in Figure 6, this chained, interactive workflow:
-  • Mesh Generation: The workflow commences with a user’s directive for a computational mesh, prompting the agent to invoke Generate_mesh to produce the geometric structure.
-  • Simulation Execution: Following successful mesh generation, the user’s instruction defines the simulation’s physical parameters such as time steps and kinematic viscosity.
-The agent receives these parameters and uses the mesh as input for Generate_simulate automatically extracted by Code2MCP to configure and execute the simulation computation, abstracting away the complex syntax of solver control files, ensuring physical parameters are correctly formatted and validated.
-This function automates the creation of complex run scripts, such as Allrun or Slurm files, and manages the execution on either a local machine or a High-Performance Computing cluster, capturing log files for subsequent review.
-  • Results Visualization: After the simulation converges, a user may request a visualization of a physical field.
-The agent then calls Visualize_velocity to render the simulation’s output.
-This function automatically processes the raw simulation data to generate insightful visual representations, such as 3D streamline plots.
-This process obviates the need for users to write complex post-processing scripts, enabling rapid analysis and interpretation of the simulation results.
-Figure 6: AI agent orchestrates a CFD simulation pipeline by sequentially invoking Generate_mesh, Generate_simulate, and Visualize_velocity functions to guide a user through the entire process, from initial mesh generation and simulation setup to the final velocity visualization.
-5 Conclusion
-Addressing the key challenge of insufficient tool supply in the MCP ecosystem, this paper introduces Code2MCP, a framework designed to automatically convert GitHub repositories into functional MCP services.
-The framework leverages a multi-agent workflow to handle core tasks of code comprehension, environment replication, and tool abstraction, while incorporating a “Run-Review-Fix” self-correction loop to enhance the reliability of the automation process.
-We demonstrate the framework’s feasibility and effectiveness by successfully converting several scientific computing libraries from fields including bioinformatics, symbolic mathematics, and computational fluid dynamics.
-This work presents an automated pathway for connecting existing codebases to the agent tool ecosystem, offering a promising solution to alleviate the MCP supply bottleneck and lower the barrier for its adoption.
-Appendix A Appendix
-A.1 Use of Large Language Models
-During manuscript preparation, large language models (LLMs) are used solely as general-purpose writing assistants for grammar checking, wording refinement, and improving clarity.
-LLMs don’t contribute to research ideation, methodological design, or experimental execution.
-All suggestions produced by LLMs are reviewed, edited, and vetted by the authors, who take full responsibility for the entire content of the paper.
-A.2 Agent Roles and System Prompts
+A.4 Agent Roles and System Prompts
 This section outlines the roles of the specialized agents within the Code2MCP framework. The following are placeholders for their system prompts.
@@
 - Summarize fragilities (optional deps, platform quirks) and propose guards (timeouts, argument validation, deterministic seeds).
-- Also produce a case description: case name, case domain, case category, case solver.
-Ensure domain/category/solver belong to case_stats[’case_domain’/’case_category’/’case_solver’].
+- Also produce a case description: case name, case domain, case category, and case solver, using a consistent taxonomy across repositories.
 - Output a compact plan for generation: candidate tools (name, brief description, inputs with types/defaults, outputs, idempotency, side effects) and environment assumptions.
@@
 - Add lightweight logging (tool name, argument schema, durations) and minimal version metadata to aid troubleshooting.
-Senior Software Engineer / Code Fixer This agent diagnoses failures and applies the smallest auditable change that restores correctness while preserving public contracts, deciding between direct fix and regeneration, using strict complete-file replacement, and avoiding multi-file edits or prose in outputs.
+Senior Software Engineer This agent diagnoses failures and applies the smallest auditable change that restores correctness while preserving public contracts, deciding between direct fix and regeneration, using strict complete-file replacement, and avoiding multi-file edits or prose in outputs.
 Review & Auto-Fix System Prompt - Triage failures: import/env, type/contract, path/I-O, dependency/version, timeout/perf, platform.
@@
 - Keep structure clear, steps verifiable, and assumptions minimal; prioritize essentials for successful adoption and integration.
-A.3 Example MCP Tool Implementations Generated by Code2MCP
+A.5 Additional Protein Case Study: ESM
+Figure 12: An AI agent processes a user’s query containing a protein sequence and mutation.
+By invoking Analyze_sequence, Predict_structure, and Predict_variant_effect functions from the generated ESM MCP server, it returns key physicochemical properties, a predicted 3D structure, and an analysis of the mutation’s effects.
+Appendix B Details of Evaluation
+B.1 GPT-4 Template Baseline Prompt
+In the GPT-4 template baseline, we use a single system prompt that instructs one GPT-4 model to generate a complete MCP service implementation for a given repository. The full prompt is given below.
+GPT-4 Template Baseline System Prompt You are a single GPT-4 model acting as an autonomous MCP wrapper generator for a Python code repository.
+You receive: - A short task description.
+- The repository README (possibly truncated).
+- A list of key source files selected by static analysis, each with its file path and content.
+- A minimal description of the MCP protocol and the FastMCP library.
+Your goal is to produce a single, self-contained Python MCP server that exposes a small set of high-value tools from this repository.
+Requirements: 1.
+Design of tools - Identify 3–8 core, stable capabilities that are useful to expose as tools.
+- For each tool, choose a concise, descriptive name and a one-sentence natural-language description.
+- Specify explicit, typed parameters (no *args or **kwargs); include default values where reasonable.
+- Define clear, JSON-serializable return values.
+Avoid returning raw Python objects that cannot be serialized.
+2.
+Implementation with FastMCP - Use the FastMCP library to implement the MCP server.
+- Implement a create_app() function that constructs and returns the MCP application.
+- Register each tool with FastMCP, binding it to the underlying repository functions or classes.
+- Prefer importing from public, documented APIs instead of private or test modules.
+- Use lazy imports for heavy optional dependencies when necessary, and return a helpful error message if a required package is missing.
+3.
+Robustness and safety - Validate all inputs (types, value ranges, file existence) before calling library functions.
+- Handle exceptions inside each tool and return a standardized error object instead of raising uncaught exceptions.
+- The standardized response format for each tool is: { "success": bool, "result": any or null, "error": string or null }.
+- Avoid network access unless the repository explicitly requires it.
+- Restrict file I/O to the workspace; do not write to absolute paths outside the project.
+4.
+Coding style - Write clean, idiomatic Python 3 code with clear function boundaries.
+- Add short docstrings for each tool describing its purpose, parameters, and return value.
+- Do not generate tests, examples, or documentation as part of the MCP server file.
+- Do not include Markdown fences such as ““‘; output only raw Python code.
+5.
+Output format - Return a single Python file containing the complete MCP server implementation.
+- Do not include explanations, comments to the user, or multiple files; only the final server code that can be saved as e.g.
+mcp_service.py.
+B.2 Evaluation Metrics and Bucketing Rules
+This section defines the evaluation metrics and repository bucketing rules used in the main experiments.
+Interaction steps.
+In the RepoMaster experiments, an interaction step is defined as a single high-level agent action that changes the external state of the system.
+Concretely, we count as one step any of the following: (1) invoking a tool (including Code2MCP-generated MCP tools and built-in RepoMaster tools), (2) editing source code files, or (3) executing the target program or tests.
+The total number of interaction steps for a task is the length of the resulting action sequence.
+Task Set Description.
+In the integration experiments with RepoMaster and OpenAgents, we use a fixed set of code-understanding and tool-usage tasks that are primarily derived from public benchmarks and representative scenarios in prior work, with a small number of additional scientific-computing cases to ensure coverage of both general and scientific settings.
+At this stage, this task set remains a relatively small, research-oriented setup rather than a fully standardized large-scale benchmark, and constructing a more systematic MCP task benchmark is left as future work.
+Task success rate and coverage.
+In the OpenAgents experiments, a task is considered successful if the agent produces a final answer that passes automatic correctness checks.
+Task success rate is the fraction of tasks in a group that are successful.
+Coverage is the fraction of tasks in a group for which the agent produces any final answer that can be checked by the automatic evaluator, regardless of whether the answer is ultimately judged correct.
+Repository buckets. To analyze the impact of dependency complexity and project style, we group repositories into four buckets based on static metadata:
+  • Low-dependency libraries: at most 10 Python dependencies inferred from requirements.txt, environment.yml, or pyproject.toml, and primarily library-style public APIs.
+  • Medium-dependency scientific libraries: between 11 and 30 Python dependencies with a library-style API and no substantial external system requirements.
+  • High-dependency libraries: more than 30 Python dependencies or additional system-level requirements (e.g., MPI stacks, GPU-accelerated frameworks, or complex external services).
+  • Script-heavy projects: projects whose primary functionality is exposed via CLI scripts, notebooks, or workflow-style pipelines rather than a stable library API; these are identified using repository structure, entry points, and README descriptions.
+Appendix C Per-repository Results
+Table 3: Repository identifiers, domains, and final outcomes.
+ID Domain Repo Final_status Main_failure_type
+B1 Biomedical facebookresearch/esm success
+B2 Biomedical biocore/scikit-bio fail env_failure
+B3 Biomedical Biopython/Biopython fail api_inference_error
+B4 Biomedical pysam-developers/pysam success
+B5 Biomedical deepchem/deepchem success
+P1 Psychology psychopy/psychopy fail untoolable_repo
+P2 Psychology pymc-devs/pymc success
+P3 Psychology sahahn/BPt success
+P4 Psychology mne-tools/mne-python fail api_inference_error
+P5 Psychology neuropsychology/NeuroKit success
+M1 Math sympy/sympy success
+M2 Math scipy/scipy fail api_inference_error
+M3 Math fredrik-johansson/mpmath success
+M4 Math cvxpy/cvxpy success
+M5 Math sagemath/sage success
+Ea1 Earth Science obspy/obspy success
+Ea2 Earth Science Unidata/MetPy fail mcp_spec_violation
+Ea3 Earth Science pyproj4/pyproj fail repo_internal_bug
+Ea4 Earth Science mapbox/rasterio success
+Ea5 Earth Science geopandas/geopandas success
+A1 Astronomy astropy/astropy success
+A2 Astronomy sunpy/sunpy fail repo_internal_bug
+A3 Astronomy lightkurve/lightkurve success
+A4 Astronomy astroML/astroML success
+A5 Astronomy astropy/astroquery success
+C1 Chemistry rdkit/rdkit success
+C2 Chemistry openbabel/openbabel fail env_failure
+C3 Chemistry bjodah/chempy success
+C4 Chemistry pyscf/pyscf fail api_inference_error
+C5 Chemistry cclib/cclib success
+Ph1 Physics csml-rpi/Foam-Agent success
+Ph2 Physics PlasmaPy/PlasmaPy fail env_failure
+Ph3 Physics pydy/pydy fail
+Ph4 Physics PyAbel/PyAbel success
+Ph5 Physics scikit-hep/scikit-hep fail untoolable_repo
+S1 Social Science networkx/networkx success
+S2 Social Science igraph/python-igraph success
+S3 Social Science networkit/networkit fail env_failure
+S4 Social Science snap-stanford/snap-python fail untoolable_repo
+S5 Social Science datamade/dedupe success
+L1 Linguistics nltk/nltk success
+L2 Linguistics explosion/spaCy success
+L3 Linguistics stanfordnlp/stanza success
+L4 Linguistics RaRe-Technologies/gensim fail import_error
+L5 Linguistics flairNLP/flair success
+E1 Econometrics statsmodels/statsmodels success
+E2 Econometrics alkaline-ml/pmdarima success
+E3 Econometrics facebook/prophet fail env_failure
+E4 Econometrics blue-yonder/tsfresh fail repo_internal_bug
+E5 Econometrics pydata/xarray success
+Table 4: Per-repository structural statistics (size, dependencies, and tests).
+ID total_files total_size Size_bucket Dependency_complexity Has_tests Has_packaging
+B1 471 33M Medium Medium Yes Yes
+B2 874 9M Medium Medium Yes Yes
+B3 1146 95M Medium High Yes Yes
+B4 561 13M Medium Medium Yes Yes
+B5 1411 136M Medium High Yes Yes
+P1 3542 65M Large High Yes Yes
+P2 371 17M Medium High Yes Yes
+P3 2669 39M Large Medium Yes Yes
+P4 1385 26M Medium High No No
+P5 496 52M Medium Medium Yes Yes
+M1 1968 29M Medium Low Yes Yes
+M2 3027 81M Large High Yes Yes
+M3 197 2M Small Low No Yes
+M4 1079 38M Medium Medium Yes Yes
+M5 5331 113M Large High Yes Yes
+Ea1 2117 34M Large High Yes Yes
+Ea2 595 99M Medium Medium Yes Yes
+Ea3 144 1M Small Medium No Yes
+Ea4 382 17M Medium Medium Yes Yes
+Ea5 351 11M Medium Medium Yes Yes
+A1 1879 24M Large High Yes Yes
+A2 877 10M Medium Medium Yes Yes
+A3 217 8M Small Medium No Yes
+A4 192 1M Small Low Yes Yes
+A5 960 23M Medium Medium Yes Yes
+C1 5740 130M Large High Yes Yes
+C2 10000 63M Large High Yes Yes
+C3 213 1M Small Low No Yes
+C4 2114 56M Medium High Yes Yes
+C5 1508 80M Large Medium Yes Yes
+Ph1 44 0.2M Small Medium Yes Yes
+Ph2 498 16M Medium Medium Yes Yes
+Ph3 263 8M Small Medium No Yes
+Ph4 243 2M Small Low Yes Yes
+Ph5 18 0.3M Small Low No Yes
+S1 951 10M Medium Low Yes Yes
+S2 248 3M Small Medium No Yes
+S3 1168 22M Medium Medium Yes Yes
+S4 678 14M Medium High Yes Yes
+S5 97 1M Small Low No No
+L1 496 8M Medium Low Yes Yes
+L2 1683 12M Medium Medium Yes Yes
+L3 574 5M Medium Medium Yes Yes
+L4 638 55M Medium Medium Yes Yes
+L5 383 5M Small Medium No Yes
+E1 2091 39M Medium High Yes Yes
+E2 225 1M Small Low No Yes
+E3 282 14M Medium Medium Yes Yes
+E4 135 10M Small Medium Yes Yes
+E5 403 8M Medium Medium Yes Yes
+Table 5: Per-repository environment and MCP conversion statistics.
+ID Env_success MCP_success Num_tools Num_tools_passed First_run_status Num_fix_rounds
+B1 Yes Yes 16 14 fail 1
+B2 No No 0 0 fail 2
+B3 Yes No 8 4 fail 2
+B4 Yes Yes 10 9 success 0
+B5 Yes Yes 14 13 fail 0
+P1 No No 0 0 fail 1
+P2 Yes Yes 9 8 success 0
+P3 Yes Yes 11 10 success 0
+P4 No No 13 12 fail 3
+P5 Yes Yes 7 7 success 0
+M1 Yes Yes 12 10 success 0
+M2 Yes No 6 2 fail 2
+M3 Yes Yes 8 7 success 0
+M4 Yes Yes 10 9 success 0
+M5 Yes Yes 15 13 fail 2
+Ea1 Yes Yes 11 8 fail 2
+Ea2 No No 5 1 fail 1
+Ea3 No No 4 0 fail 2
+Ea4 Yes Yes 9 8 success 0
+Ea5 Yes Yes 7 7 success 1
+A1 Yes Yes 8 5 fail 2
+A2 No No 3 1 fail 1
+A3 Yes Yes 10 9 success 0
+A4 Yes Yes 6 6 success 1
+A5 Yes Yes 8 8 success 0
+C1 Yes Yes 18 16 fail 1
+C2 No No 0 0 fail 2
+C3 Yes Yes 7 7 success 0
+C4 Yes No 12 10 fail 3
+C5 Yes Yes 9 8 success 1
+Ph1 Yes Yes 8 7 success 1
+Ph2 No No 0 0 fail 2
+Ph3 Yes No 6 6 fail 0
+Ph4 Yes Yes 5 5 success 0
+Ph5 No No 0 0 fail 1
+S1 Yes Yes 10 9 success 0
+S2 Yes Yes 8 7 fail 1
+S3 No No 0 0 fail 2
+S4 Yes No 0 0 fail 1
+S5 Yes Yes 5 5 success 0
+L1 Yes Yes 7 7 success 0
+L2 Yes Yes 11 10 fail 2
+L3 Yes Yes 9 8 success 1
+L4 Yes No 0 0 fail 3
+L5 Yes Yes 6 6 success 0
+E1 Yes Yes 13 11 fail 1
+E2 Yes Yes 7 7 success 0
+E3 No No 0 0 fail 2
+E4 No No 0 0 fail 1
+E5 Yes Yes 8 7 success 1
+Appendix D Example MCP Tool Implementations Generated by Code2MCP
 To illustrate the concrete, high-quality output of Code2MCP, this section presents several MCP tool implementations that were autonomously generated by Code2MCP.
 These examples are drawn from the ESM and SymPy case studies discussed in the main paper and demonstrate the framework’s ability to produce clean, robust, and immediately usable code.
-A.3.1 Tools Generated from the ESM Repository
+D.1 Tools Generated from the ESM Repository
 MCP Tool for Protein Sequence Analysis @mcp.tool(name="analyze_sequence", description="Analyze protein sequence features.") def analyze_sequence(sequence: str): """Analyzes physicochemical properties of a protein sequence.""" try: import re try: from esm import analysis except Exception: try: from .esm import analysis except Exception: import types analysis = types.SimpleNamespace( calculate_molecular_weight=lambda s: float(len(s)) * 110.0, calculate_isoelectric_point=lambda s: 7.0, calculate_instability_index=lambda s: 40.0 ) aa_set = set("ACDEFGHIKLMNPQRSTVWY") seq = re.sub(r"[^A-Za-z]", "", sequence or "").upper() seq = "".join([c for c in seq if c in aa_set]) length = len(seq) composition = {$a$: seq.count($a$) for $a$ in aa_set if seq.count($a$) > 0} molecular_weight = analysis.calculate_molecular_weight(seq) isoelectric_point = analysis.calculate_isoelectric_point(seq) instability_index = analysis.calculate_instability_index(seq) properties = { "length": length, "composition": composition, "molecular_weight": molecular_weight, "isoelectric_point": isoelectric_point, "instability_index": instability_index, "sequence": seq, } return {"success": True, "result": properties, "error": None} except Exception as e: return {"success": False, "result": None, "error": f"Error during sequence analysis: {str(e)}"}
 […]
-A.3.2 Tools Generated from the SymPy Repository
+D.1.1 Tools Generated from the SymPy Repository
 MCP Tool for Solving Equations @mcp.tool(name="solve_equation") def solve_equation(equation: str, variable: str): """ Solve equation for variable """ try: from sympy import sympify, symbols, solve, Basic def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x expr = sympify(equation) var = symbols(variable) res = solve(expr, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
 MCP Tool for Solving Linear Systems @mcp.tool(name="solve_linear_system") def solve_linear_system(system: list, variables: list): """ Solve system of linear equations """ try: from sympy import sympify, symbols, linsolve, Basic def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x eqs = [sympify(e) for e in system] vars_sym = [symbols(v) for v in variables] res = linsolve(eqs, vars_sym) return {"success": True, "result": ser(list(res))} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Differentiation @mcp.tool(name="differentiate") def differentiate(expr: str, variable: str): """ Calculate derivative of expression with respect to variable """ try: from sympy import sympify, symbols, diff, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = diff(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Integration @mcp.tool(name="integrate_expression") def integrate_expression(expr: str, variable: str): """ Calculate integral of expression with respect to variable """ try: from sympy import sympify, symbols, integrate, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = integrate(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Differentiation @mcp.tool(name="differentiate") def differentiate(expr: str, variable: str): """ Calculate derivative of expression """ try: from sympy import sympify, symbols, diff, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = diff(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Integration @mcp.tool(name="integrate_expression") def integrate_expression(expr: str, variable: str): """ Calculate integral of expression """ try: from sympy import sympify, symbols, integrate, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = integrate(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
 MCP Tool for Polynomial Creation @mcp.tool(name="create_polynomial") def create_polynomial(expr: str, variable: str = None): """ Create polynomial from expression """ try: from sympy import sympify, symbols, Poly, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) if variable: var = symbols(variable) res = Poly(expr_sym, var) else: res = Poly(expr_sym) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
 MCP Tool for Polynomial Factoring @mcp.tool(name="factor_polynomial") def factor_polynomial(poly: str): """ Factor polynomial expression """ try: from sympy import sympify, factor, Basic def ser(x): if isinstance(x, Basic): return str(x) return x poly_sym = sympify(poly) res = factor(poly_sym) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Fourier Transform @mcp.tool(name="fourier_transform") def fourier_transform(expression: str, time_var: str, freq_var: str): """ Calculate Fourier transform """ try: from calc.symbolic import sympify, symbols import numpy as np from scipy import integrate expr = sympify(expression) t = symbols(time_var) omega = symbols(freq_var) def integrand(t_val): return float(expr.subs(t, t_val)) t_range = (-50, 50) omega_vals = np.linspace(-10, 10, 100) result = [] for omega_val in omega_vals: def f(t_val): return integrand(t_val) * np.exp(-1j * omega_val * t_val) integral_result, _ = scipy_integrate.quad(f, t_range[0], t_range[1], limit=100) result.append(complex(integral_result)) return {"success": True, "result": [str(complex(r)) for r in result]} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Laplace Transform @mcp.tool(name="laplace_transform") def laplace_transform(expression: str, time_var: str, laplace_var: str): """ Calculate Laplace transform of a function.
-""" try: from calc.symbolic import sympify, symbols from calc.transforms import laplace_transform as sympy_laplace_transform import numpy as np from scipy import integrate expr = sympify(expression) t = symbols(time_var) s = symbols(laplace_var) try: F = sympy_laplace_transform(expr, t, s, noconds=True) return {"success": True, "result": str(F)} except Exception: def integrand(t_val): return float(expr.subs(t, t_val)) t_range = (0, 100) s_vals = np.linspace(0.1, 10, 50) result = [] for s_val in s_vals: def f(t_val): return integrand(t_val) * np.exp(-s_val * t_val) integral_result, _ = scipy_integrate.quad(f, t_range[0], t_range[1], limit=100) result.append({"s": float(s_val), "F": str(complex(integral_result))}) return {"success": True, "result": result} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Z Transform @mcp.tool(name="z_transform") def z_transform(expression: str, time_var: str, z_var: str, limit: int = 100): """ Calculate Z transform """ try: def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x expr = sympify(expression) n = symbols(time_var) z = symbols(z_var) try: result = Sum(expr * (z ** (-n)), (n, 0, $o$)).doit() result = simplify(result) return {"success": True, "result": ser(result)} except Exception: try: if expr.is_Pow and expr.base.is_symbol and expr.exp == n: a = expr.base result = z / (z - a) return {"success": True, "result": ser(result)} elif expr.is_Mul and len(expr.args) == 2: for arg in expr.args: if arg == n: other = expr / n if other.is_Pow and other.base.is_symbol and other.exp == n: a = other.base result = a * z / (z - a)**2 return {"success": True, "result": ser(result)} elif expr.is_Mul and n**2 in expr.args: other = expr / n**2 if other.is_Pow and other.base.is_symbol and other.exp == n: a = other.base result = a * z * (z + a) / (z - a)**3 return {"success": True, "result": ser(result)} result = 0 for k in range(limit + 1): term = expr.subs(n, k) * (z ** (-k)) result += term return {"success": True, "result": ser(result)} except Exception: result = 0 for k in range(limit + 1): term = expr.subs(n, k) * (z ** (-k)) result += term return {"success": True, "result": ser(result)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-A.3.3 Tools Generated from the Foam-Agent Repository
+MCP Tool for Fourier Transform from sympy import sympify, symbols, fourier_transform as sympy_fourier_transform, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="fourier_transform") def fourier_transform_tool(expression: str, time_var: str = "t", freq_var: str = "w"): try: expr = sympify(expression) t = symbols(time_var) omega = symbols(freq_var) F = sympy_fourier_transform(expr, t, omega) return {"success": True, "result": _serialize(F), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Laplace Transform from sympy import sympify, symbols, laplace_transform as sympy_laplace_transform, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="laplace_transform") def laplace_transform_tool(expression: str, time_var: str = "t", laplace_var: str = "s"): try: expr = sympify(expression) t = symbols(time_var) s = symbols(laplace_var) F, _, _ = sympy_laplace_transform(expr, t, s) return {"success": True, "result": _serialize(F), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Z Transform from sympy import sympify, symbols, summation, $o$, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="z_transform") def z_transform_tool(expression: str, time_var: str = "n", z_var: str = "z", limit: int = 100): try: expr = sympify(expression) n = symbols(time_var) z = symbols(z_var) try: result = summation(expr * z**(-n), (n, 0, $o$)) except Exception: result = summation(expr * z**(-n), (n, 0, limit)) return {"success": True, "result": _serialize(result), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+D.1.2 Tools Generated from the Foam-Agent Repository
 MCP Tool for Mesh Generation @mcp.tool(name="generate_mesh", description="Generate computational mesh using Foam-Agent internals.") def generate_mesh(requirements: str, case_dir: str = "./output", mesh_mode: str = "gmsh", custom_mesh_path: str | None = None): try: from src.config import Config from src.main import initialize_state from src.nodes.meshing_node import meshing_node config = Config() config.case_dir = case_dir state = initialize_state(user_requirement=requirements, config=config, custom_mesh_path=custom_mesh_path) if mesh_mode == "custom": state["mesh_type"] = "custom_mesh" elif mesh_mode == "gmsh": state["mesh_type"] = "gmsh_mesh" else: state["mesh_type"] = "standard_mesh" res = meshing_node(state) return {"success": True, "result": res, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}

### v4

@@
-Transforming Code Repositories into MCP Services
-Abstract
+Code2MCP: Transforming Code Repositories into MCP Services
+Abstract.
 The Model Context Protocol (MCP) aims to create a standard for how Large Language Models use tools.
@@
 By providing a novel automated pathway to unlock GitHub, the world’s largest code repository, for the MCP ecosystem, Code2MCP serves as a catalyst to significantly accelerate the protocol’s adoption and practical application.
-The code is public at https://github.com/DEFENSE-SEU/Code2MCP.
-1 Introduction
-The landscape of artificial intelligence is increasingly defined by autonomous agents that leverage Large Language Models (LLMs) to interact with external tools (Wang et al., 2024; bubeck2023sparks).
+The code is public at https://anonymous.4open.science/r/code2mcp-7E34.
+Agentic AI, Multi-Agent, Model Context Protocol, Large Language Model, GitHub
+copyright: none
+1. Introduction
+The landscape of artificial intelligence is increasingly defined by autonomous agents that leverage Large Language Models (LLMs) to interact with external tools (Wang et al., 2024).
 To overcome the inherent limitations of LLMs in tasks requiring real-time information or precise computation, the paradigm of tool-augmented reasoning has become central (Huang et al., 2024; Hao et al., 2023).
-Seminal works have demonstrated that models can effectively learn to invoke external functions (Schick et al., 2023; Qin et al., 2023; Parisi et al., 2022).
-However, this burgeoning ecosystem faces a fundamental scalability challenge: the $N\times M$ integration problem (Li et al., 2023; Liang et al., 2023; Qin et al., 2023; Anthropic, 2023).
+Seminal works have demonstrated that models can effectively learn to invoke external functions (Schick et al., 2023; Qin et al., 2024b; Parisi et al., 2022).
+For example, Toolformer (Schick et al., 2023) shows that LLMs can be trained to invoke external APIs in a self-supervised manner, establishing the basic feasibility of tool use.
+Later work further scales this paradigm by orchestrating larger tool and model collections, such as Gorilla (Patil et al., 2024) and HuggingGPT (Shen et al., 2023), showing that tool use can be extended beyond a small, fixed set of predefined APIs.
+However, this burgeoning ecosystem faces a fundamental scalability challenge: the $N\times M$ integration problem (Liang et al., 2024; Qin et al., 2024b; Anthropic, 2023).
 Each of the N models or agent applications often requires a bespoke connector for each of the M tools it must access.
@@
 To address this, MCP is proposed as a universal standard that specifies how agents and tools should communicate, enabling an interoperable “plug-and-play” ecosystem (Anthropic, 2023).
-In response to this integration challenge, the community’s efforts have evolved, inadvertently revealing a deeper, more foundational bottleneck (Yue et al., 2025).
-The initial challenge is to establish the fundamental feasibility of tool use, where a limited set of tools proves the feasibility of the paradigm (Schick et al., 2023; Patil et al., 2023; Ding et al., 2025; Qin et al., 2024).
+Figure 1.
+While most research focuses on the consumption of tools (right side), one bottleneck is their supply (left side).
+Code2MCP solves the supply problem by converting the code repository into a standardized MCP-compliant tool.
+Table 1. A comparative summary of related works in tool-augmented LLMs.
+Type Work Core Contribution Tool Pool Tool Selection MCP
+Consumer Toolformer (Schick et al., 2023) Teaching LLM to use external tools 5 predefined tools Fine-tuning $\times$
+SciToolAgent (Ding et al., 2025) Domain-specific enhancement for scientific tool utilization KG of Scientific Tools (500+ tools) Retrieval on KG $\times$
+HuggingGPT (Shen et al., 2023) Increase the size of tool pool huggingface.co LLM task planning $\times$
+Gorilla (Patil et al., 2024) TorchHub, TensorHub (1600+ tools) Retriever-aware training $\times$
+OpenAgents (Xie et al., 2023) Tool use from open-source beyond a closed pool github.com Multi-agent planning $\times$
+RepoMaster (Wang and others, 2025) Rule-based deep search $\times$
+RAG-MCP (Gan and Sun, 2025) RAG for tool selection from MCP MCP.so Retrieval on MCP $\checkmark$
+Supplier Code2MCP GitHub Repo to standardized MCP github.com N/A $\checkmark$
+In response to this integration challenge, the community’s efforts have evolved, inadvertently revealing a deeper, more foundational bottleneck (Yue et al., 2025a).
+The initial challenge is to establish the fundamental feasibility of tool use, where a limited set of tools proves the feasibility of the paradigm (Schick et al., 2023; Patil et al., 2024; Ding et al., 2025; Qin et al., 2024a).
 To break past the inherent scarcity of these platforms, the focus shifts to the vast landscape of open-source repositories (Wang and others, 2025; Xie et al., 2023; Wang, 2023).
@@
 The MCP emerges as a direct answer to this chaos, promising a universal interface (Yu et al., 2025; Anthropic, 2023).
-Yet, this leads to the critical gap: research is focused almost exclusively on the consumption side of MCP, using services from a presumed-to-exist pool (Gan and Sun, 2025), while the foundational “supply-side” problem of how to populate this pool from existing software is largely unaddressed.
-Figure 1: While most research focuses on the consumption of tools (right side), one bottleneck is their supply (left side).
-Code2MCP solves the supply problem by converting the code repository into a standardized MCP-compliant tool.
-However, while these efforts advance the consumption side of the problem, how agents can better use tools, they largely overlook a more fundamental bottleneck on the supply side.
-This supply bottleneck is not a theoretical concern but a stark reality preventing standards like MCP from achieving widespread adoption.
-For example, RAG-MCP (Gan and Sun, 2025) utilizes over 4,400 servers on mcp.so, but there are 268 million public GitHub repositories.
+Yet, this leads to a critical gap: existing research focuses almost exclusively on the consumption side of MCP, using services from a presumed-to-exist pool (Gan and Sun, 2025), while the foundational supply-side problem of how to populate this pool from existing software remains largely unaddressed.
+We identify and formalize this largely overlooked tool supply problem in the MCP ecosystem: how to automatically transform existing code repositories into reusable, registrable, and agent-consumable MCP services, rather than relying on one-off, repository-level execution.
+As illustrated in Figure 1, existing approaches make different trade-offs among the large, open, and standard dimensions of tool pools, while the key bottleneck of the MCP pool lies in the insufficient supply of standardized tool services.
+Code2MCP is designed to act as a supplier that fills this gap.
+However, while these efforts advance the consumption side of the problem by improving how agents use tools, they largely overlook a more fundamental bottleneck on the supply side.
+This bottleneck is not theoretical.
+Most consumer-side settings assume that a usable tool pool already exists, whether it is a curated knowledge graph as in SciToolAgent (Ding et al., 2025) or open-source repositories as explored by OpenAgents (Xie et al., 2023) and RepoMaster (Wang and others, 2025).
+Yet the scale of available standardized tools remains limited.
+For example, RAG-MCP (Gan and Sun, 2025) performs retrieval over a pre-existing MCP server registry, with mcp.so as a representative platform, whereas the broader software ecosystem contains hundreds of millions of public GitHub repositories.
 The critical question of how to create a large and diverse pool of these standardized, agent-ready tools has been left unaddressed.
 This creates a major adoption gap, effectively locking away the largest software repository, GitHub, from this emerging ecosystem.
+Figure 2. Overview of the Code2MCP framework. The system takes a GitHub repository URL as input and automatically generates a complete MCP service through a multi-agent workflow.
 In this paper, we introduce Code2MCP, a new framework designed to bridge the critical tool supply gap.
-Code2MCP presents a blueprint for transforming any GitHub repository into a functional and documented MCP service with minimal human intervention.
+Code2MCP presents a blueprint for transforming a wide range of GitHub repositories into functional and documented MCP services with minimal human intervention.
 However, this transformation is a complex endeavor encompassing four pivotal challenges: (1) deep code comprehension to identify core functionalities, (2) reliable environment replication to ensure executability, (3) intelligent tool abstraction to design useful and valid service interfaces, and (4) robust self-correction to handle the inevitable errors throughout the process.
-To systematically address these challenges, as shown in Figure 2, Code2MCP implements a collaborative multi-agent system (Park et al., 2023).
-Unlike general-purpose coding agents (Cognition, 2024; Jimenez et al., 2024; Zhang et al., 2024b; Huang et al., 2023), different agents in our framework are specialized for the distinct stages of code analysis, environment setup, and API design.
-Crucially, the overall reliability of this workflow is ensured by an integrated Run-Review-Fix self-correction cycle, which endows the system with the ability to autonomously debug and refine the entire conversion process.
+To systematically address these challenges, as shown in Figure 2, Code2MCP implements a collaborative multi-agent system (Park et al., 2023; Yue et al., 2024; Tran et al., 2025; Guo et al., 2024; Dorri et al., 2018).
+Unlike general-purpose coding agents (Cognition, 2024; Jimenez et al., 2024; Zhang et al., 2024), different agents in our framework are specialized for the distinct stages of code analysis, environment setup, and API design.
+Crucially, the overall reliability of this workflow is ensured by an integrated Run-Review-Fix self-correction cycle, where failure traces are explicitly analyzed to generate targeted modification plans that constrain subsequent regeneration to specific files and code regions.
+This design differs from naive retry-based generation by enforcing localized, diagnosis-driven repairs rather than global re-synthesis, while keeping previously validated tool interfaces stable whenever possible.
 The key contributions of this work are listed as follows:
@@
 This provides a concrete and practical pathway to enrich the MCP ecosystem with specialized, high-value tools.
-Table 1: A comparative summary of related works in tool-augmented LLMs.
-Work Core Contribution Tool Pool Tool Selection MCP
-Consumer Toolformer Teaching LLM to use external tools 5 predefined tools Fine-tuning $\times$
-SciToolAgent Domain-specific enhancement for scientific tool utilization KG of Scientific Tools (500+ tools) Retrieval on KG $\times$
-HuggingGPT Increase the size of tool pool huggingface.co LLM task planning $\times$
-Gorilla TorchHub, TensorHub (1600+ tools) Retriever-aware training $\times$
-OpenAgents Tool use from open-source beyond a closed pool github.com Multi-agent planning $\times$
-RepoMaster Rule-based deep search $\times$
-RAG-MCP RAG for tool selection from MCP MCP.so (4,400+) Retrieval on MCP $\checkmark$
-Supplier Code2MCP GitHub Repo to standardized MCP github.com N/A $\checkmark$
-2 Related Work
+2. Related Work
 As summarized in Section 1 and Table 1, the pioneering works focus on progressively expanding the scope of tool use, from initial feasibility studies using a few predefined APIs to leveraging large, curated tool platforms, and ultimately, to the ambitious goal of directly interfacing with unstructured open-source repositories.
@@
 These approaches significantly expand the number of available tools.
-For instance, Gorilla (Patil et al., 2023) fine-tunes models on a massive corpus of API calls from hubs like TorchHub and TensorFlow Hub.
+For instance, Gorilla (Patil et al., 2024) fine-tunes models on a massive corpus of API calls from hubs like TorchHub and TensorFlow Hub.
 Similarly, HuggingGPT (Shen et al., 2023) positions an LLM as a controller to delegate tasks to specialized models within the Hugging Face ecosystem.
@@
 These works confront the complexity of real-world code but expose the core bottleneck: the vast majority of these repositories are not designed for programmatic use by LLM agents.
-They lack standardized interfaces (Zhang et al., 2024a; Jin et al., 2024; Ray, 2025), forcing the agent into an ad-hoc, brittle, and unreliable process of reverse-engineering the code, setting up its environment, and managing dependencies for every single task.
-This chaotic integration process illustrates a critical failure on the tool supply side.
+They lack standardized interfaces (Zhang et al., 2025; Jin et al., 2024; Ray, 2025), forcing the agent into an ad-hoc, brittle, and unreliable process of reverse-engineering the code, setting up its environment, and managing dependencies for every single task.
+While these systems partially touch the supply side by enabling agents to operate over open-source repositories, they do not aim to produce standardized, reusable MCP services that can be shared across agents and tasks.
 The Emergence of Standardization and Unaddressed Gap.
 Recognizing this chaos, the community has moved towards standardization, exemplified by the MCP.
-For example, RAG-MCP (Gan and Sun, 2025) explores how an agent can effectively retrieve and select the most appropriate MCP service from a pool of 4400+ available options.
+For example, RAG-MCP (Gan and Sun, 2025) explores how an agent can effectively retrieve and select the most appropriate MCP service from a pool of available options.
 This approach is promising, but it presumes the existence of a rich ecosystem of MCP-compliant services (Hasan et al., 2025).
-This highlights a crucial, unaddressed gap: how is this ecosystem of MCP services populated in the first place?
-Figure 2: Overview of the Code2MCP framework. The system takes a GitHub repository URL as input and automatically generates a complete MCP service through a multi-agent workflow.
-3 Methodology: The Code2MCP Framework
-To achieve the goal of automatically transforming an arbitrary GitHub repository into a fully functional and reliable MCP service, we design Code2MCP, an automated framework driven by the collaboration of seven specialized agents.
+This highlights a crucial gap: how is this ecosystem of MCP services populated in the first place?
+3. Methodology: The Code2MCP Framework
+To automatically transform an arbitrary GitHub repository into a fully functional and reliable MCP service, we design Code2MCP, an automated framework driven by the collaboration of seven specialized agents.
 The entire conversion process, as depicted in Figure 2, is a multi-agent workflow that begins with code analysis, proceeds through a core Run-Review-Fix self-correction loop, and culminates in the generation of a merge-ready pull request.
-For completeness, a more formal description of this workflow is provided in the appendix A.3.
+A detailed description of the framework is provided in Appendix 1.
 Suppose there exists a consumer-side work listed in Table 1 that finds a suitable GitHub repository that may solve the user’s query.
@@
 This Run-Review-Fix loop repeats until tests pass or a maximum of $B$ attempts is reached.
+Figure 3.
+The complete output directory structure generated by the Code2MCP framework.
+The top-level workspace contains the original repository alongside all generated artifacts within the mcp_output directory.
 Finalization and Delivery.
@@
 To facilitate review and adoption by the original repository maintainers, it also generates a README file explaining how to use the new MCP service.
-All artifacts are arranged into a reproducible directory structure under the workspace (Appendix A.2), and the agent can prepare a pull request to submit these additions back to the original repository.
-4 Evaluation
-4.1 Experimental Setup
+As shown in Figure 3, All artifacts are arranged into a reproducible directory structure under the workspace, and the agent can prepare a pull request to submit these additions back to the original repository.
+Threat model and safety boundaries.
+Code2MCP executes third-party code during environment setup and validation, which poses inherent security risks.
+In our evaluation, all conversions are run inside isolated worker environments with restricted filesystem scope and bounded runtime budgets.
+By default, outbound network access is disabled during validation; for repositories whose core functionality intrinsically depends on external services, network access is selectively enabled under an explicit allowlist and treated as a separate evaluation condition.
+We do not attempt to systematically evaluate adversarial repositories or malicious payloads in this work; instead, our security discussion focuses on operational safeguards during automated conversion rather than adversarial robustness.
+4. Evaluation
+4.1. Experimental Setup
 Task.
-The evaluation task is: given a GitHub repository URL, Code2MCP automatically generates an MCP service for that repository, and the outcome of the conversion is judged as a success or a failure.
+The evaluation task is: given a GitHub repository, Code2MCP automatically generates an MCP service for that repository, and the outcome of the conversion is judged as a success or a failure.
 The evaluation focuses on the overall conversion success rate across different domains and repository characteristics, and on the time and stability of Code2MCP compared with manual implementation and a GPT-4 template-based baseline on the same set of repositories.
@@
 Code2MCP is evaluated on 50 open-source repositories from 10 scientific and engineering domains (5 per domain), covering biomedical science, psychology, mathematical computing, earth science, chemistry, physics, astronomy, social science, linguistics, and econometrics.
-These repositories range from small utility libraries to large frameworks with complex dependency stacks.
+The goal of this selection is not exhaustive coverage, but to expose the framework to a diverse range of dependency structures, API styles, and execution environments commonly encountered in scientific software.
 For each repository, the evaluation logs the environment construction result, MCP conversion result, number of generated tools, and failure types observed across all Run-Review-Fix rounds; per-repository details and the associated failure labels are summarized in a large table in the Appendix C, and a subset of representative scientific-computing repositories is further used for the Run-Review-Fix ablation and qualitative case studies.
 Success criterion.
-A conversion is regarded as successful only if the following three conditions all hold: the runtime environment can be reconstructed according to the dependency specifications provided by the repository, the generated MCP server can be started and must pass a unified RunNode test, and at least three core tools can be invoked correctly in automatic tests and return valid outputs.
-If any of these conditions is not satisfied, the repository is counted as a failure.
-Failure taxonomy. To analyze failure modes, all failed repositories are annotated with one or more labels from the following six categories:
+A conversion is counted as successful iff all of the following hold: (1) the runtime environment can be reconstructed from repository-provided specifications; (2) the generated MCP server starts and passes the unified RunNode validation; (3) at least three pre-identified, documentation-facing functionalities succeed under automatic invocation.
+These functionalities are selected deterministically from the repository’s public documentation and matched to generated tools by name and semantic description, prior to conversion.
+Failure taxonomy.
+Each failed repository is annotated with one or more failure labels derived from execution artifacts (installation logs, server logs, and schema validation).
+If multiple error types occur across Run–Review–Fix rounds, labels are aggregated; a primary label is also recorded based on the final failing stage.
   • env_failure. The environment or dependencies cannot be reconstructed.
@@
 Case studies are conducted on servers equipped with 8 NVIDIA H100 80 GB GPUs.
-Table 2: Per-domain summary of environment setup success, basic test success, recovery by the Run-Review-Fix (RRF) loop (“–” if none are recovered), and final MCP conversion success.
-Domain Repos Env succ. Test succ. RRF recovered Avg rounds MCP succ.
+Table 2. Per-domain summary of environment setup success, basic test success, recovery by the Run-Review-Fix (RRF) loop (“–” if none are recovered), and final MCP conversion success.
+Domain Repos Env success Test success RRF recovered Avg rounds MCP success
 Biomedical 5 4/5 (80%) 2/5 (40%) 1/2 (50%) 1.5 3/5 (60%)
@@
 Overall 50 38/50 (76%) 24/50 (48%) 8/19 (42.1%) 1.4 32/50 (64%)
-4.2 Large-scale Repository Conversion and Failure Analysis
-Overall and per-domain success.
-The overall conversion success rate is first analyzed at the domain level.
+4.2. Large-scale Repository Conversion and Failure Analysis
+Overall and per-domain success. The overall conversion success rate is first analyzed at the domain level.
+(a) Failure label distribution.
+(b) RRF ablation results.
+Figure 4.
+(a) summarizes the distribution of 42 failure labels assigned to 18 failed repositories; a single repository can trigger multiple failure types across different Run-Review-Fix rounds.
+(b) reports repository success rates with and without Run‑Review‑Fix and the remaining errors across RRF rounds for five representative scientific MCP repositories.
 In Table 2, for each domain, the number of repositories with successful environment setup and basic test passing, the number of initially failing repositories recovered by the Run-Review-Fix loop and their average rounds, and the final number of successful MCP conversions.
@@
 Per-repository results, including environment construction and MCP conversion outcomes as well as the number of generated tools, are summarized in the large table in the appendix and can be further grouped by domain or repository size if desired.
-Figure 3: Distribution of 42 failure labels assigned to the 18 failed repositories. A single repository can trigger multiple failure types across different Run-Review-Fix rounds.
 Failure modes.
@@
 Labels are aggregated across all Run-Review-Fix rounds, so a single repository may receive multiple labels if different error types are observed during different stages of the pipeline.
-The distribution of these labels is visualized in Figure 3.
+The distribution of these labels is visualized in Figure 4(a).
 The labels env_failure and api_inference_error account for more than half of all failures, indicating that environment reconstruction and interface inference are the primary bottlenecks of the current workflow.
 The labels import_error and untoolable_repo are also relatively common, reflecting the difficulty of handling complex import paths, initialization order, and script-driven repositories in a fully automated pipeline.
+Figure 5.
+Task success rates (top row) and completion time (bottom row) for Human experts, GPT-4 with a template, and Code2MCP across representative repositories in four repository groups.
+The x-axis lists representative repositories in each group, labeled by domain abbreviation and index (e.g., Math-1).
 Run-Review-Fix ablation.
 To assess the specific contribution of the Run-Review-Fix self-correction loop, an ablation study is conducted on a set of representative scientific-computing repositories by comparing two configurations: a single-pass configuration without the loop and the full Code2MCP pipeline with the loop enabled.
-Figure 4 summarizes the results.
+Figure 4(b) summarizes the results.
 The bar chart on the left shows that the per-repository success rates increase for most repositories once the loop is enabled, often moving from medium success under a single-generation setting to success rates close to or on par with human-written wrappers.
 The line chart on the right plots the average number of remaining errors as a function of the number of Run-Review-Fix rounds, where “remaining errors” denote the number of failing assertions or uncaught exceptions in the unified test after each round.
-These results indicate that the majority of initial failures can be automatically repaired within one to three Run-Review-Fix rounds, whereas the few repositories that still fail after multiple rounds represent the main current limitation of the method.
-4.3 Comparison with Human and GPT-4 Baselines
+These results indicate the majority of initial failures can be automatically repaired within one to three Run-Review-Fix rounds, whereas the few repositories that still fail after multiple rounds represent the main limitation of the method.
+4.3. Comparison with Human and GPT-4 Baselines
 The evaluation compares three configurations on the same set of representative repositories:
@@
   • the Code2MCP configuration, which uses the multi-agent pipeline with the Run-Review-Fix loop.
-Figure 4: Repo success rates with and without Run‑Review‑Fix (left) and remaining errors across RRF rounds (right) for five representative scientific MCP repositories.
 In the human configuration, ten graduate students with at least three years of Python experience and basic familiarity with MCP concepts implement wrappers from scratch.
 Before the study, all participants read the MCP specification and a commented example repository.
-Each participant is randomly assigned several repositories, and the measured time is from the moment they start reading the repository to the point where the minimal test passes.
-In the GPT-4 template configuration, a unified system prompt instructs a single GPT-4 model to generate a complete MCP service implementation based on the repository README and several key source files selected by static analysis.
-We use the same base model and decoding configuration as in Code2MCP, namely gpt-4o-2024-05-13 with temperature set to 1.0, to ensure a fair comparison.
-The model is allowed up to three dialogue turns: the first turn generates an initial implementation, and up to two additional turns may revise the code based on execution errors returned by a test runner.
-No additional agents or planning mechanisms are used, and the full system prompt is provided in Appendix B.1.
-Figure 5: Average task success rate (left) and average completion time (right) across the ten scientific domains for the three configurations: Human experts, GPT-4 with basic scaffolding, and Code2MCP.
+Each participant is randomly assigned several repositories.
+For consistency, the same success criterion and the same validation procedure are applied.
+Completion time is measured from the start of wrapper implementation to the first time the unified validation passes, or to termination under the same budgets.
+In the GPT-4 template configuration, a unified system prompt instructs a single model to generate a complete MCP service implementation based on the repository README and several key source files selected by static analysis.
+The model performs end-to-end interface design, code generation, execution, and iterative debugging based on execution tracebacks, constituting a single-agent baseline.
+The baseline follows the same validation procedure and the same success criterion.
+Generation and revision are performed iteratively based on execution errors returned by the test runner, and the process stops when the unified validation passes or when the evaluation budget is exhausted.
+No additional agents, retrieval modules, or external planning tools are used.
+The full system prompt is provided in Appendix E.1.
+(a) RepoMaster with/without Code2MCP.
+(b) OpenAgents with/without Code2MCP.
+Figure 6.
+(a) shows RepoMaster with and without Code2MCP tools, reporting task success rates and average interaction steps across task groups.
+(b) shows OpenAgents with and without Code2MCP tools, reporting task success rates and coverage across general and scientific task groups.
 In the Code2MCP configuration, the default multi-agent pipeline with the Run-Review-Fix loop is used.
-For each repository, we record the final success status and the completion time, defined as the end-to-end wall-clock time starting from the beginning of environment setup and ending when the unified MCP test first passes or the run is terminated; see Appendix B.2 for details.
+For each repository, we record the final success status and the completion time, defined as the end-to-end wall-clock time starting from the beginning of environment setup and ending when the unified MCP test first passes or the run is terminated.
+To make the evaluation protocol fully self-contained.
 Figure 7 summarizes the average task success rates and completion times of the three configurations across the ten scientific domains.
@@
 Completion time is the average end-to-end wall-clock time per attempt, measured from the beginning of environment setup until the minimal MCP test first passes or the run is terminated.
-Figure 6: Task success rates of Human experts, GPT-4 with a template, and Code2MCP on representative repositories in the four repository groups.
-The x-axis lists representative repositories in each group, labeled by domain abbreviation and index (e.g., Math-1).
-The overall trend is as follows: Code2MCP achieves task success rates close to human implementations in most domains, and consistently higher than the GPT-4 template configuration.
-In terms of completion time, human implementations are typically on the order of hours, while Code2MCP completes in tens of minutes; the GPT-4 template configuration is the fastest but has lower success rates.
-When repositories are grouped by dependency complexity and project type into low-dependency libraries, medium-dependency scientific libraries, high-dependency libraries, and script-heavy projects, Figures 6 and 7 further break down the same comparison on representative repositories within each group.
-The bucketing rules are detailed in Appendix B.2.
-Figure 7: Completion time of Human experts, GPT-4 with a template, and Code2MCP on representative repositories in the four repository groups.
-4.4 Integration with Existing Tool Systems
+The overall trend is as follows: Code2MCP achieves task success rates close to human implementations in most domains, and higher than the GPT-4 template configuration.
+In terms of completion time, human implementations are on the order of hours, while Code2MCP completes in tens of minutes; the GPT-4 template configuration is the fastest but has lower success rates.
+When repositories are grouped by dependency complexity and project type into low-dependency libraries, medium-dependency scientific libraries, high-dependency libraries, and script-heavy projects, Figures 5 further break down the same comparison on representative repositories within each group.
+4.4. Integration with Existing Tool Systems
 This subsection summarizes how Code2MCP integrates with existing tool ecosystems by combining it with RepoMaster and OpenAgents on a shared set of tool-usage tasks.
+Task Set Description.
+In the integration experiments with RepoMaster and OpenAgents, we evaluate on a fixed set of code-understanding and tool-usage tasks designed to assess whether automatically generated MCP tools can be discovered, selected, and invoked by existing agent systems under realistic usage patterns.
+The task categories and evaluation protocol are kept identical across methods, and full task specifications and examples are provided in Appendix D.
 RepoMaster + Code2MCP.
-In the RepoMaster setting, code-understanding and tool-usage tasks are run in two configurations: using RepoMaster alone to operate at the source-code level, and using RepoMaster together with Code2MCP, where RepoMaster first selects relevant repositories and then invokes the corresponding MCP tools.
-Figure 8 shows that the configuration with Code2MCP achieves higher task success rates and requires fewer interaction steps (see Appendix B.2 for metric definitions), suggesting that delegating part of the work to standardized MCP tools improves efficiency without modifying the upstream agent policy.
+In the RepoMaster setting, tool-usage and code-understanding tasks are run in two configurations: using RepoMaster alone to operate at the source-code level, and using RepoMaster together with Code2MCP, where RepoMaster first selects relevant repositories and then invokes the corresponding MCP tools.
+Figure 6(a) shows that the configuration with Code2MCP achieves higher task success rates and requires fewer interaction steps, suggesting that delegating part of the work to standardized MCP tools improves efficiency.
+Figure 7.
+Average task success rate (left) and average completion time (right) across the ten scientific domains for the three configurations: Human experts, GPT-4 with basic scaffolding, and Code2MCP.
 OpenAgents + Code2MCP.
 In the OpenAgents setting, cross-domain tasks are evaluated under two configurations: using OpenAgents alone, and using OpenAgents with MCP tools automatically generated by Code2MCP added to the tool pool.
-As shown in Figure 9, we report the task success rate and coverage, where coverage and success rate are defined in Appendix B.2.
+A task is considered successful if the agent produces a final answer that passes automatic correctness checks.
+(a) ESM protein analysis workflow.
+(b) Foam-Agent CFD workflow.
+Figure 8.
+(a) shows a protein analysis pipeline where an agent invokes Analyze_sequence, Predict_structure, and Predict_variant_effect from the generated ESM MCP server to return physicochemical properties, a predicted 3D structure, and mutation impact analysis.
+(b) shows a CFD pipeline where the agent calls Generate_mesh, Generate_simulate, and Visualize_velocity to guide the user from mesh generation and solver setup through simulation execution and final velocity-field visualization.
+Task success rate is the fraction of tasks in a group that are successful.
+Coverage is the fraction of tasks in a group for which the agent produces any final answer that can be checked by the automatic evaluator, regardless of whether the answer is ultimately judged correct.
+As shown in Figure 6(b), we report the task success rate and coverage.
 Adding Code2MCP tools increases both metrics across all task groups, indicating that a richer MCP tool pool can be effectively exploited by existing planning and retrieval strategies.
 Overall, RepoMaster and OpenAgents handle tool discovery and code retrieval on the consumer side, while Code2MCP supplies additional MCP services on the supply side.
-Under the same front-end selection and planning strategies, integrating Code2MCP enriches the tool pool and yields measurable improvements in end-to-end task performance.
-Figure 8: RepoMaster with and without Code2MCP tools: task success rate (left) and average interaction steps (right) across task groups.
-4.5 Case Studies: Protein, Math, and Computational Fluid Dynamics
-This subsection highlights three representative repositories from the 50-repository evaluation set: the biomedical protein modeling library ESM, the symbolic mathematics library SymPy, and the CFD framework Foam-Agent built on OpenFOAM, illustrating the kinds of MCP tools Code2MCP generates and how agents use them in practice.
-Figure 9: OpenAgents with and without Code2MCP tools: task success rate (left) and coverage (right) across general and scientific task groups.
+Under the same front-end selection and planning strategies, integrating Code2MCP enriches the tool pool and yields measurable improvements.
+4.5. Case Studies: Protein, Math, and Computational Fluid Dynamics
+This subsection highlights three representative repositories from the 50-repository evaluation set: the biomedical protein modeling library ESM, the symbolic mathematics library SymPy, and the CFD framework Foam-Agent (Yue et al., 2025b) built on OpenFOAM, illustrating the kinds of MCP tools Code2MCP generates and how agents use them in practice.
 Bridging protein science with AI agents.
-In protein science, models such as AlphaFold (Jumper et al., 2021) greatly improve structure prediction, but using them in everyday research typically requires substantial scripting and environment setup.
-ESM models complement AlphaFold with efficient sequence modeling and zero-shot variant effect prediction, yet traditionally still require multi-step Python pipelines for computing physicochemical properties, predicting structures, and analyzing mutations.
+In protein science, models such as AlphaFold (Jumper et al., 2021) have greatly improved structure prediction, but using them in everyday research still often requires substantial scripting and environment setup.
+ESM models complement AlphaFold with efficient sequence modeling and zero-shot variant effect prediction, yet many common analyses remain multi-step Python pipelines for computing physicochemical properties, predicting structures, and interpreting mutations.
+This workflow can be distracting for researchers who primarily focus on wet-lab experiments and want results that are easy to reproduce and explain.
 Code2MCP converts ESM into an MCP service by exposing tools such as Analyze_sequence, Predict_structure, and Predict_variant_effect.
-We provide a qualitative example of using ESM-based MCP tools for protein analysis in Appendix A.5.
-Figure 10: An AI agent orchestrates a CFD simulation pipeline by invoking Generate_mesh, Generate_simulate, and Visualize_velocity functions to guide a user through the entire process, from initial mesh generation and simulation setup to the final velocity visualization.
+Figure 8(a) illustrates a qualitative example.
+When a user provides a protein sequence and a mutation site like “Q145G”, the agent can first call Analyze_sequence to produce a baseline characterization, including molecular weight, amino acid composition, and hydrophobicity.
+It then calls Predict_structure to generate a 3D structure and uses a Predicted Aligned Error (PAE) plot to assess reliability.
+Finally, Predict_variant_effect helps locate the mutation on the structure and summarize likely impacts, and it can also generate a simple visualization to clarify the mutation’s spatial context.
 Enhancing mathematical reliability in AI agents.
@@
 Code2MCP organizes common capabilities from SymPy into a mathematical MCP service with tools for limits and integration, matrix operations, symbolic simplification, and transforms such as Fourier and Laplace.
-When asked to compute an integral or the volume of a solid of revolution, an agent simply invokes the corresponding tool and returns an exact answer, relying on SymPy as a backend instead of ad-hoc prompt-based reasoning.
+For example, to find the indefinite integral of $y=\ln(x^{2}+1)$, it calls the corresponding integration tool, and for the volume of revolution of $y=\ln(x)\cdot\sin(x)$ from $x=1$ to $x=2\pi$, it invokes the volume tool to produce the exact result.
+By grounding sensitive mathematical steps in validated code rather than prompt-based manipulation, the workflow becomes easier to verify and reproduce.
 Automating CFD simulation for AI agents.
-CFD workflows typically involve mesh generation, solver configuration, execution, and post-processing, traditionally coordinated by engineers through shell scripts and configuration files.
-Foam-Agent provides a higher-level interface to OpenFOAM, and Code2MCP further turns it into an MCP service with tools such as Generate_mesh, Generate_simulate, and Visualize_velocity.
-As illustrated in Figure 10, a user describes the target geometry and physical parameters in natural language, and the agent calls these tools in sequence to construct the mesh, run the simulation, and visualize the velocity field.
-5 Conclusion
-Addressing the key challenge of insufficient tool supply in the MCP ecosystem, this paper introduces Code2MCP, a framework that automatically converts GitHub repositories into functional MCP services.
-The framework employs a multi-agent workflow for code comprehension, environment reconstruction, and tool abstraction, augmented by a “Run-Review-Fix” self-correction loop that improves end-to-end reliability across diverse scientific libraries.
-Our evaluation on 50 repositories shows that Code2MCP can robustly expose high-value APIs as MCP tools while also revealing clear failure modes and boundary conditions.
-This work presents an automated pathway for connecting existing codebases to the agent tool ecosystem, and we leave systematically hardening security and building standardized large-scale MCP benchmarks from these converted services as important directions for future work.
+CFD workflows typically involve mesh generation, solver configuration, execution, and visualization, traditionally coordinated by engineers through shell scripts and configuration files—an error-prone and labor-intensive process.
+Foam-Agent provides a higher-level interface to OpenFOAM, and Code2MCP further turns it into an MCP service with tools such as Generate_mesh, Generate_simulate, and Visualize_velocity, allowing an agent to run the workflow step-by-step in dialogue.
+As illustrated in Figure 8(b), a user describes the target geometry and physical parameters in natural language, and the agent calls these tools in sequence to construct the mesh, run the simulation, and visualize the velocity field.
+5. Conclusion
+This paper addresses the key challenge of insufficient tool supply in the MCP ecosystem by introducing Code2MCP, a framework that automatically converts GitHub repositories into functional MCP services.
+The framework employs a multi-agent workflow for code comprehension, environment reconstruction, and tool abstraction, augmented by a “Run-Review-Fix” self-correcting loop that improves end-to-end reliability across diverse scientific libraries.
+An evaluation on 50 repositories demonstrates that Code2MCP can robustly expose high-value APIs as MCP tools while also revealing clear failure modes and boundary conditions.
+Overall, the results highlight an automated pathway for connecting existing codebases to the agent tool ecosystem, while surfacing concrete limitations that delimit the feasibility of fully automated tool supply.
 Appendix A Appendix
-A.1 Use of Large Language Models
+A.1. Use of Large Language Models
 During manuscript preparation, large language models (LLMs) are used solely as general-purpose writing assistants for grammar checking, wording refinement, and improving clarity.
@@
 All suggestions produced by LLMs are reviewed, edited, and vetted by the authors, who take full responsibility for the entire content of the paper.
-A.2 Output Directory Structure
-Figure 11 shows the final directory layout generated by Code2MCP, where the original repository and the synthesized MCP artifacts are organized under a unified workspace.
-Figure 11: The complete output directory structure generated by the Code2MCP framework.
-The top-level workspace contains the original repository alongside all generated artifacts within the mcp_output directory.
-A.3 Detailed Conversion Pipeline and Algorithm
-This section outlines the roles of the specialized agents within the Code2MCP framework and their coordination as shown in Algorithm 1, where Download and Environment initialize the workspace, Analysis and Generation construct the MCP service, Run and Review form the Run-Review-Fix loop, and Finalize produces the final deliverables.
-The following are placeholders for their system prompts.
+A.2. algorithm
 Algorithm 1 The Code2MCP Framework
@@
 18:Output: A merge-ready Pull Request containing the functional MCP service.
-A.4 Agent Roles and System Prompts
-This section outlines the roles of the specialized agents within the Code2MCP framework. The following are placeholders for their system prompts.
-Environment Agent This agent rapidly provisions a minimal, isolated runtime for the repository, with minimal dependencies and a short smoke test; if setup fails, propose one lightweight, auditable fallback without modifying the repository; keep all steps reproducible and pragmatic.
-Environment System Prompt - Prefer Conda; use venv only if Conda is unavailable or clearly unsuitable.
-- Detect dependency sources by priority: environment.yml > requirements.txt > pyproject.toml > setup files; never guess hidden dependencies.
-- Pin versions when explicit; otherwise install the minimal viable set.
-Prefer CPU wheels unless GPU is explicitly required.
-- Normalize cross-platform behavior; avoid absolute paths; use relative POSIX-like paths; ensure UTF-8 locale.
-- Smoke test: print Python version and platform; import fastmcp; attempt to import the project’s top-level package or a primary CLI; exit code 0 indicates pass.
-- On failure, capture exact command, exit code, last 80 lines of stderr, and timing; propose exactly one minimal remedy (e.g., switch to venv, install a single missing package, try one version pin, extend timeout once).
-- Apply at most one fallback; never change repository code; do not write outside the workspace; do not weaken security (e.g., no SSL bypass).
-- Cache wheels where possible; avoid global pollution; record reproducible commands and resolved versions.
-- Default to offline validation; if network is strictly required, justify briefly and bound the scope.
-- Emit a compact environment report (platform, Python, manager, explicit deps, resolved pins, pass/fail).
-Code Analysis Expert This agent performs static analysis to shape the repository into a compact, high-value capability surface, selecting stable public functionality, filtering out test/demo code, and producing a concise plan aligned with predefined domains, categories, and solvers.
-Code Analysis System Prompt - Ingest repository signals (structure, import graph, README/docstrings, CLI entry points; DeepWiki if available) to identify stable public APIs suitable as MCP tools.
-- Prefer documented, side-effect-bounded surfaces; exclude tests, internals, notebooks, long demos unless clearly valuable and controllable.
-- Define crisp tool boundaries: explicit inputs/outputs, preconditions/postconditions, resource needs (CPU/GPU/memory/time), and I/O constraints.
-- Note minimal adapter needs (path normalization, dtype coercion, lazy imports) and hazards (network access, file mutation, global state).
-- Summarize fragilities (optional deps, platform quirks) and propose guards (timeouts, argument validation, deterministic seeds).
-- Also produce a case description: case name, case domain, case category, and case solver, using a consistent taxonomy across repositories.
-- Output a compact plan for generation: candidate tools (name, brief description, inputs with types/defaults, outputs, idempotency, side effects) and environment assumptions.
-Keep it actionable and minimal.
-Code Generation Expert This agent synthesizes a robust MCP service from the analysis plan with clean design, consistent interfaces, graceful failure handling, and immediate executability, defining clear tool endpoints, enforcing explicit typed parameters and standardized returns, and avoiding test or example tooling.
-Code Generation System Prompt - Produce clean, runnable Python (no Markdown fences).
-Use FastMCP to build the MCP service.
-- Implement create_app() that returns the service; register tools with concise names and user-facing descriptions.
-- For every tool: explicit, typed parameters (no *args/**kwargs); validate inputs; JSON-serializable outputs.
-- Standard return shape: success: bool, result: any or null, error: string or null.
-- Handle optional dependencies via lazy imports; emit helpful errors without crashing the service; prefer CPU fallbacks when reasonable.
-- Ensure deterministic defaults (fixed seeds when relevant); avoid hidden global state; restrict file I/O to the workspace with existence/size/extension checks.
-- Design for cross-platform paths; avoid shell-specific behavior; bound execution time and memory.
-- Do not generate tests as tools.
-Expose a small set of high-value, composable endpoints; avoid overexposing internals.
-- Add lightweight logging (tool name, argument schema, durations) and minimal version metadata to aid troubleshooting.
-Senior Software Engineer This agent diagnoses failures and applies the smallest auditable change that restores correctness while preserving public contracts, deciding between direct fix and regeneration, using strict complete-file replacement, and avoiding multi-file edits or prose in outputs.
-Review & Auto-Fix System Prompt - Triage failures: import/env, type/contract, path/I-O, dependency/version, timeout/perf, platform.
-- Choose minimal direct fix vs.
-regeneration; provide a one-line rationale and confidence (low/med/high).
-Prefer adapter-boundary mitigations (lazy import, existence checks, parameter coercion).
-- Apply strict complete-file replacement for the single target file; return pure code only; do not alter unrelated sections.
-- Preserve interface contracts and standardized error shapes; add narrow guards instead of broad catches.
-- Enforce cross-platform path handling and deterministic behavior; do not introduce external network calls or new side effects.
-- Optionally add a tiny internal sanity check if it prevents recurrence without bloat.
-- Bound attempts (<= B).
-If still failing, emit a compact escalation note: failing step, last command, stderr tail, and the next best single remediation.
-Final Agent This agent consolidates artifacts and workflow logs into developer-facing documentation and delivery notes that are precise, reproducible, and integration-ready, producing a concise README with installation, quick start, key tools, troubleshooting guidance, and references.
-Final Agent System Prompt - Write a concise developer README (Markdown) including: 1) Overview and value; roles of MCP and FastMCP; supported OS.
-2) Minimal reproducible install (Conda/venv), commands, pinned dependencies, offline notes; Windows PowerShell and Linux shell variants.
-3) Quick start to launch the service and call 2–3 key tools with copy-pasteable snippets and basic error handling.
-4) Tool list: endpoint name, one-line description, key parameters (types/defaults), return shape, idempotency/side effects, typical runtime class.
-5) Troubleshooting: environment/import issues, optional deps, timeouts, path problems, permissions, CPU/GPU enablement; any bounded network caveats.
-6) Reproducibility and telemetry: how to capture environment report, versions, minimal repro commands; where logs/artifacts live.
-7) License and compliance notes: repository license, usage constraints, safety guardrails.
-- Keep structure clear, steps verifiable, and assumptions minimal; prioritize essentials for successful adoption and integration.
-A.5 Additional Protein Case Study: ESM
-Figure 12: An AI agent processes a user’s query containing a protein sequence and mutation.
-By invoking Analyze_sequence, Predict_structure, and Predict_variant_effect functions from the generated ESM MCP server, it returns key physicochemical properties, a predicted 3D structure, and an analysis of the mutation’s effects.
-Appendix B Details of Evaluation
-B.1 GPT-4 Template Baseline Prompt
-In the GPT-4 template baseline, we use a single system prompt that instructs one GPT-4 model to generate a complete MCP service implementation for a given repository. The full prompt is given below.
-GPT-4 Template Baseline System Prompt You are a single GPT-4 model acting as an autonomous MCP wrapper generator for a Python code repository.
-You receive: - A short task description.
-- The repository README (possibly truncated).
-- A list of key source files selected by static analysis, each with its file path and content.
-- A minimal description of the MCP protocol and the FastMCP library.
-Your goal is to produce a single, self-contained Python MCP server that exposes a small set of high-value tools from this repository.
-Requirements: 1.
-Design of tools - Identify 3–8 core, stable capabilities that are useful to expose as tools.
-- For each tool, choose a concise, descriptive name and a one-sentence natural-language description.
-- Specify explicit, typed parameters (no *args or **kwargs); include default values where reasonable.
-- Define clear, JSON-serializable return values.
-Avoid returning raw Python objects that cannot be serialized.
-2.
-Implementation with FastMCP - Use the FastMCP library to implement the MCP server.
-- Implement a create_app() function that constructs and returns the MCP application.
-- Register each tool with FastMCP, binding it to the underlying repository functions or classes.
-- Prefer importing from public, documented APIs instead of private or test modules.
-- Use lazy imports for heavy optional dependencies when necessary, and return a helpful error message if a required package is missing.
-3.
-Robustness and safety - Validate all inputs (types, value ranges, file existence) before calling library functions.
-- Handle exceptions inside each tool and return a standardized error object instead of raising uncaught exceptions.
-- The standardized response format for each tool is: { "success": bool, "result": any or null, "error": string or null }.
-- Avoid network access unless the repository explicitly requires it.
-- Restrict file I/O to the workspace; do not write to absolute paths outside the project.
-4.
-Coding style - Write clean, idiomatic Python 3 code with clear function boundaries.
-- Add short docstrings for each tool describing its purpose, parameters, and return value.
-- Do not generate tests, examples, or documentation as part of the MCP server file.
-- Do not include Markdown fences such as ““‘; output only raw Python code.
-5.
-Output format - Return a single Python file containing the complete MCP server implementation.
-- Do not include explanations, comments to the user, or multiple files; only the final server code that can be saved as e.g.
-mcp_service.py.
-B.2 Evaluation Metrics and Bucketing Rules
-This section defines the evaluation metrics and repository bucketing rules used in the main experiments.
-Interaction steps.
-In the RepoMaster experiments, an interaction step is defined as a single high-level agent action that changes the external state of the system.
-Concretely, we count as one step any of the following: (1) invoking a tool (including Code2MCP-generated MCP tools and built-in RepoMaster tools), (2) editing source code files, or (3) executing the target program or tests.
-The total number of interaction steps for a task is the length of the resulting action sequence.
-Task Set Description.
-In the integration experiments with RepoMaster and OpenAgents, we use a fixed set of code-understanding and tool-usage tasks that are primarily derived from public benchmarks and representative scenarios in prior work, with a small number of additional scientific-computing cases to ensure coverage of both general and scientific settings.
-At this stage, this task set remains a relatively small, research-oriented setup rather than a fully standardized large-scale benchmark, and constructing a more systematic MCP task benchmark is left as future work.
-Task success rate and coverage.
-In the OpenAgents experiments, a task is considered successful if the agent produces a final answer that passes automatic correctness checks.
-Task success rate is the fraction of tasks in a group that are successful.
-Coverage is the fraction of tasks in a group for which the agent produces any final answer that can be checked by the automatic evaluator, regardless of whether the answer is ultimately judged correct.
-Repository buckets. To analyze the impact of dependency complexity and project style, we group repositories into four buckets based on static metadata:
-  • Low-dependency libraries: at most 10 Python dependencies inferred from requirements.txt, environment.yml, or pyproject.toml, and primarily library-style public APIs.
-  • Medium-dependency scientific libraries: between 11 and 30 Python dependencies with a library-style API and no substantial external system requirements.
-  • High-dependency libraries: more than 30 Python dependencies or additional system-level requirements (e.g., MPI stacks, GPU-accelerated frameworks, or complex external services).
-  • Script-heavy projects: projects whose primary functionality is exposed via CLI scripts, notebooks, or workflow-style pipelines rather than a stable library API; these are identified using repository structure, entry points, and README descriptions.
+Appendix B Example MCP Tool Implementations Generated by Code2MCP
+To illustrate the concrete, high-quality output of Code2MCP, this section presents several MCP tool implementations that were autonomously generated by Code2MCP.
+These examples are drawn from the ESM and SymPy case studies discussed in the main paper and demonstrate the framework’s ability to produce clean, robust, and immediately usable code.
+B.1. Tools Generated from the ESM Repository
+MCP Tool for Protein Sequence Analysis @mcp.tool(name="analyze_sequence", description="Analyze protein sequence features (validated, no placeholder fallback).") def analyze_sequence(sequence: str): """ Analyze physicochemical properties of a protein sequence.
+Returns real computed values; never uses placeholder fallbacks.
+""" import re import time start = time.time() try: aa_set = set("ACDEFGHIKLMNPQRSTVWY") seq = re.sub(r"[^A-Za-z]", "", sequence or "").upper() seq = "".join([c for c in seq if c in aa_set]) if not seq: return {"success": False, "result": None, "error": "Empty/invalid protein sequence after filtering."} composition = {$a$: seq.count($a$) for $a$ in sorted(aa_set) if seq.count($a$) > 0} try: from Bio.SeqUtils.ProtParam import ProteinAnalysis # type: ignore except Exception as e: return { "success": False, "result": None, "error": ( "Biopython is required for physicochemical analysis but is not installed.
+" "Install: pip install biopython.
+" f"Original import error: {str(e)}" ), } pa = ProteinAnalysis(seq) molecular_weight = float(pa.molecular_weight()) isoelectric_point = float(pa.isoelectric_point()) instability_index = float(pa.instability_index()) aromaticity = float(pa.aromaticity()) gravy = float(pa.gravy()) # hydrophobicity helix, turn, sheet = pa.secondary_structure_fraction() sec_struct = {"helix": float(helix), "turn": float(turn), "sheet": float(sheet)} duration_ms = int((time.time() - start) * 1000) result = { "sequence": seq, "length": len(seq), "composition": composition, "molecular_weight": molecular_weight, "isoelectric_point": isoelectric_point, "instability_index": instability_index, "aromaticity": aromaticity, "gravy": gravy, "secondary_structure_fraction": sec_struct, "metadata": {"backend": "biopython_protparam", "duration_ms": duration_ms}, } return {"success": True, "result": result, "error": None} except Exception as e: return {"success": False, "result": None, "error": f"Error during sequence analysis: {str(e)}"}
+[…]
+B.1.1. Tools Generated from the SymPy Repository
+MCP Tool for Solving Equations @mcp.tool(name="solve_equation") def solve_equation(equation: str, variable: str): """ Solve equation for variable """ try: from sympy import sympify, symbols, solve, Basic def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x expr = sympify(equation) var = symbols(variable) res = solve(expr, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Solving Linear Systems @mcp.tool(name="solve_linear_system") def solve_linear_system(system: list, variables: list): """ Solve system of linear equations """ try: from sympy import sympify, symbols, linsolve, Basic def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x eqs = [sympify(e) for e in system] vars_sym = [symbols(v) for v in variables] res = linsolve(eqs, vars_sym) return {"success": True, "result": ser(list(res))} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Differentiation @mcp.tool(name="differentiate") def differentiate(expr: str, variable: str): """ Calculate derivative of expression """ try: from sympy import sympify, symbols, diff, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = diff(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Integration @mcp.tool(name="integrate_expression") def integrate_expression(expr: str, variable: str): """ Calculate integral of expression """ try: from sympy import sympify, symbols, integrate, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = integrate(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Polynomial Creation @mcp.tool(name="create_polynomial") def create_polynomial(expr: str, variable: str = None): """ Create polynomial from expression """ try: from sympy import sympify, symbols, Poly, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) if variable: var = symbols(variable) res = Poly(expr_sym, var) else: res = Poly(expr_sym) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Polynomial Factoring @mcp.tool(name="factor_polynomial") def factor_polynomial(poly: str): """ Factor polynomial expression """ try: from sympy import sympify, factor, Basic def ser(x): if isinstance(x, Basic): return str(x) return x poly_sym = sympify(poly) res = factor(poly_sym) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Fourier Transform from sympy import sympify, symbols, fourier_transform as sympy_fourier_transform, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="fourier_transform") def fourier_transform_tool(expression: str, time_var: str = "t", freq_var: str = "w"): try: expr = sympify(expression) t = symbols(time_var) omega = symbols(freq_var) F = sympy_fourier_transform(expr, t, omega) return {"success": True, "result": _serialize(F), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Laplace Transform from sympy import sympify, symbols, laplace_transform as sympy_laplace_transform, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="laplace_transform") def laplace_transform_tool(expression: str, time_var: str = "t", laplace_var: str = "s"): try: expr = sympify(expression) t = symbols(time_var) s = symbols(laplace_var) F, _, _ = sympy_laplace_transform(expr, t, s) return {"success": True, "result": _serialize(F), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Z Transform from sympy import sympify, symbols, summation, $o$, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="z_transform") def z_transform_tool(expression: str, time_var: str = "n", z_var: str = "z", limit: int = 100): try: expr = sympify(expression) n = symbols(time_var) z = symbols(z_var) try: result = summation(expr * z**(-n), (n, 0, $o$)) except Exception: result = summation(expr * z**(-n), (n, 0, limit)) return {"success": True, "result": _serialize(result), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+B.1.2. Tools Generated from the Foam-Agent Repository
+MCP Tool for Mesh Generation @mcp.tool(name="generate_mesh", description="Generate computational mesh using Foam-Agent internals.") def generate_mesh(requirements: str, case_dir: str = "./output", mesh_mode: str = "gmsh", custom_mesh_path: str | None = None): try: from src.config import Config from src.main import initialize_state from src.nodes.meshing_node import meshing_node config = Config() config.case_dir = case_dir state = initialize_state(user_requirement=requirements, config=config, custom_mesh_path=custom_mesh_path) if mesh_mode == "custom": state["mesh_type"] = "custom_mesh" elif mesh_mode == "gmsh": state["mesh_type"] = "gmsh_mesh" else: state["mesh_type"] = "standard_mesh" res = meshing_node(state) return {"success": True, "result": res, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Simulation Generation and Run @mcp.tool(name="generate_simulate", description="Write inputs and run simulation via Foam-Agent graph.") def generate_simulate(requirements: str, case_dir: str = "./output", custom_mesh_path: str | None = None, run_mode: str = "auto"): try: from src.config import Config from src.main import create_foam_agent_graph, initialize_state config = Config() config.case_dir = case_dir state = initialize_state(user_requirement=requirements, config=config, custom_mesh_path=custom_mesh_path) if custom_mesh_path: state["mesh_type"] = "custom_mesh" if run_mode == "local": state["cluster_info"] = None elif run_mode == "hpc": state["cluster_info"] = {"scheduler": "slurm"} workflow = create_foam_agent_graph().compile() workflow.invoke(state) return {"success": True, "result": {"case_dir": config.case_dir}, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+MCP Tool for Velocity Visualization @mcp.tool(name="visualize_velocity", description="Post-process and visualize velocity (|U|, streamlines, slices).") def visualize_velocity(case_dir: str, plot_type: str = "magnitude", plane: str | None = "xy"): try: from src.config import Config from src.main import initialize_state from src.nodes.visualization_node import visualization_node config = Config() config.case_dir = case_dir state = initialize_state(user_requirement="", config=config, custom_mesh_path=None) state["case_dir"] = case_dir state["visualization_request"] = {"plot_type": plot_type, "plane": plane} vis_res = visualization_node(state) return {"success": True, "result": vis_res, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
 Appendix C Per-repository Results
-Table 3: Repository identifiers, domains, and final outcomes.
+Table 3. Repository identifiers, domains, and final outcomes.
   ID   Domain   Repo   Final_status   Main_failure_type
-B1 Biomedical facebookresearch/esm success
+  B1   Biomedical   facebookresearch/esm   success   –
   B2   Biomedical   biocore/scikit-bio   fail   env_failure
   B3   Biomedical   Biopython/Biopython   fail   api_inference_error
-B4 Biomedical pysam-developers/pysam success
-B5 Biomedical deepchem/deepchem success
+  B4   Biomedical   pysam-developers/pysam   success   –
+  B5   Biomedical   deepchem/deepchem   success   –
   P1   Psychology   psychopy/psychopy   fail   untoolable_repo
-P2 Psychology pymc-devs/pymc success
-P3 Psychology sahahn/BPt success
+  P2   Psychology   pymc-devs/pymc   success   –
+  P3   Psychology   sahahn/BPt   success   –
   P4   Psychology   mne-tools/mne-python   fail   api_inference_error
-P5 Psychology neuropsychology/NeuroKit success
-M1 Math sympy/sympy success
+  P5   Psychology   neuropsychology/NeuroKit   success   –
+  M1   Math   sympy/sympy   success   –
   M2   Math   scipy/scipy   fail   api_inference_error
-M3 Math fredrik-johansson/mpmath success
-M4 Math cvxpy/cvxpy success
-M5 Math sagemath/sage success
-Ea1 Earth Science obspy/obspy success
+  M3   Math   fredrik-johansson/mpmath   success   –
+  M4   Math   cvxpy/cvxpy   success   –
+  M5   Math   sagemath/sage   success   –
+  Ea1   Earth Science   obspy/obspy   success   –
   Ea2   Earth Science   Unidata/MetPy   fail   mcp_spec_violation
   Ea3   Earth Science   pyproj4/pyproj   fail   repo_internal_bug
-Ea4 Earth Science mapbox/rasterio success
-Ea5 Earth Science geopandas/geopandas success
-A1 Astronomy astropy/astropy success
+  Ea4   Earth Science   mapbox/rasterio   success   –
+  Ea5   Earth Science   geopandas/geopandas   success   –
+  A1   Astronomy   astropy/astropy   success   –
   A2   Astronomy   sunpy/sunpy   fail   repo_internal_bug
-A3 Astronomy lightkurve/lightkurve success
-A4 Astronomy astroML/astroML success
-A5 Astronomy astropy/astroquery success
-C1 Chemistry rdkit/rdkit success
+  A3   Astronomy   lightkurve/lightkurve   success   –
+  A4   Astronomy   astroML/astroML   success   –
+  A5   Astronomy   astropy/astroquery   success   –
+  C1   Chemistry   rdkit/rdkit   success   –
   C2   Chemistry   openbabel/openbabel   fail   env_failure
-C3 Chemistry bjodah/chempy success
+  C3   Chemistry   bjodah/chempy   success   –
   C4   Chemistry   pyscf/pyscf   fail   api_inference_error
-C5 Chemistry cclib/cclib success
-Ph1 Physics csml-rpi/Foam-Agent success
+  C5   Chemistry   cclib/cclib   success   –
+  Ph1   Physics   csml-rpi/Foam-Agent   success   –
   Ph2   Physics   PlasmaPy/PlasmaPy   fail   env_failure
-Ph3 Physics pydy/pydy fail
-Ph4 Physics PyAbel/PyAbel success
+  Ph3   Physics   pydy/pydy   fail   –
+  Ph4   Physics   PyAbel/PyAbel   success   –
   Ph5   Physics   scikit-hep/scikit-hep   fail   untoolable_repo
-S1 Social Science networkx/networkx success
-S2 Social Science igraph/python-igraph success
+  S1   Social Science   networkx/networkx   success   –
+  S2   Social Science   igraph/python-igraph   success   –
   S3   Social Science   networkit/networkit   fail   env_failure
   S4   Social Science   snap-stanford/snap-python   fail   untoolable_repo
-S5 Social Science datamade/dedupe success
-L1 Linguistics nltk/nltk success
-L2 Linguistics explosion/spaCy success
-L3 Linguistics stanfordnlp/stanza success
+  S5   Social Science   datamade/dedupe   success   –
+  L1   Linguistics   nltk/nltk   success   –
+  L2   Linguistics   explosion/spaCy   success   –
+  L3   Linguistics   stanfordnlp/stanza   success   –
   L4   Linguistics   RaRe-Technologies/gensim   fail   import_error
-L5 Linguistics flairNLP/flair success
-E1 Econometrics statsmodels/statsmodels success
-E2 Econometrics alkaline-ml/pmdarima success
+  L5   Linguistics   flairNLP/flair   success   –
+  E1   Econometrics   statsmodels/statsmodels   success   –
+  E2   Econometrics   alkaline-ml/pmdarima   success   –
   E3   Econometrics   facebook/prophet   fail   env_failure
   E4   Econometrics   blue-yonder/tsfresh   fail   repo_internal_bug
-E5 Econometrics pydata/xarray success
-Table 4: Per-repository structural statistics (size, dependencies, and tests).
+  E5   Econometrics   pydata/xarray   success   –
+Table 4. Per-repository structural statistics (size, dependencies, and tests).
 ID total_files total_size Size_bucket Dependency_complexity Has_tests Has_packaging
@@
 E5 403 8M Medium Medium Yes Yes
-Table 5: Per-repository environment and MCP conversion statistics.
+Table 5. Per-repository environment and MCP conversion statistics.
 ID Env_success MCP_success Num_tools Num_tools_passed First_run_status Num_fix_rounds
@@
 E5 Yes Yes 8 7 success 1
-Appendix D Example MCP Tool Implementations Generated by Code2MCP
-To illustrate the concrete, high-quality output of Code2MCP, this section presents several MCP tool implementations that were autonomously generated by Code2MCP.
-These examples are drawn from the ESM and SymPy case studies discussed in the main paper and demonstrate the framework’s ability to produce clean, robust, and immediately usable code.
-D.1 Tools Generated from the ESM Repository
-MCP Tool for Protein Sequence Analysis @mcp.tool(name="analyze_sequence", description="Analyze protein sequence features.") def analyze_sequence(sequence: str): """Analyzes physicochemical properties of a protein sequence.""" try: import re try: from esm import analysis except Exception: try: from .esm import analysis except Exception: import types analysis = types.SimpleNamespace( calculate_molecular_weight=lambda s: float(len(s)) * 110.0, calculate_isoelectric_point=lambda s: 7.0, calculate_instability_index=lambda s: 40.0 ) aa_set = set("ACDEFGHIKLMNPQRSTVWY") seq = re.sub(r"[^A-Za-z]", "", sequence or "").upper() seq = "".join([c for c in seq if c in aa_set]) length = len(seq) composition = {$a$: seq.count($a$) for $a$ in aa_set if seq.count($a$) > 0} molecular_weight = analysis.calculate_molecular_weight(seq) isoelectric_point = analysis.calculate_isoelectric_point(seq) instability_index = analysis.calculate_instability_index(seq) properties = { "length": length, "composition": composition, "molecular_weight": molecular_weight, "isoelectric_point": isoelectric_point, "instability_index": instability_index, "sequence": seq, } return {"success": True, "result": properties, "error": None} except Exception as e: return {"success": False, "result": None, "error": f"Error during sequence analysis: {str(e)}"}
-[…]
-D.1.1 Tools Generated from the SymPy Repository
-MCP Tool for Solving Equations @mcp.tool(name="solve_equation") def solve_equation(equation: str, variable: str): """ Solve equation for variable """ try: from sympy import sympify, symbols, solve, Basic def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x expr = sympify(equation) var = symbols(variable) res = solve(expr, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Solving Linear Systems @mcp.tool(name="solve_linear_system") def solve_linear_system(system: list, variables: list): """ Solve system of linear equations """ try: from sympy import sympify, symbols, linsolve, Basic def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x eqs = [sympify(e) for e in system] vars_sym = [symbols(v) for v in variables] res = linsolve(eqs, vars_sym) return {"success": True, "result": ser(list(res))} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Differentiation @mcp.tool(name="differentiate") def differentiate(expr: str, variable: str): """ Calculate derivative of expression """ try: from sympy import sympify, symbols, diff, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = diff(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Integration @mcp.tool(name="integrate_expression") def integrate_expression(expr: str, variable: str): """ Calculate integral of expression """ try: from sympy import sympify, symbols, integrate, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = integrate(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Polynomial Creation @mcp.tool(name="create_polynomial") def create_polynomial(expr: str, variable: str = None): """ Create polynomial from expression """ try: from sympy import sympify, symbols, Poly, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) if variable: var = symbols(variable) res = Poly(expr_sym, var) else: res = Poly(expr_sym) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Polynomial Factoring @mcp.tool(name="factor_polynomial") def factor_polynomial(poly: str): """ Factor polynomial expression """ try: from sympy import sympify, factor, Basic def ser(x): if isinstance(x, Basic): return str(x) return x poly_sym = sympify(poly) res = factor(poly_sym) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Fourier Transform from sympy import sympify, symbols, fourier_transform as sympy_fourier_transform, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="fourier_transform") def fourier_transform_tool(expression: str, time_var: str = "t", freq_var: str = "w"): try: expr = sympify(expression) t = symbols(time_var) omega = symbols(freq_var) F = sympy_fourier_transform(expr, t, omega) return {"success": True, "result": _serialize(F), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Laplace Transform from sympy import sympify, symbols, laplace_transform as sympy_laplace_transform, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="laplace_transform") def laplace_transform_tool(expression: str, time_var: str = "t", laplace_var: str = "s"): try: expr = sympify(expression) t = symbols(time_var) s = symbols(laplace_var) F, _, _ = sympy_laplace_transform(expr, t, s) return {"success": True, "result": _serialize(F), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Z Transform from sympy import sympify, symbols, summation, $o$, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="z_transform") def z_transform_tool(expression: str, time_var: str = "n", z_var: str = "z", limit: int = 100): try: expr = sympify(expression) n = symbols(time_var) z = symbols(z_var) try: result = summation(expr * z**(-n), (n, 0, $o$)) except Exception: result = summation(expr * z**(-n), (n, 0, limit)) return {"success": True, "result": _serialize(result), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-D.1.2 Tools Generated from the Foam-Agent Repository
-MCP Tool for Mesh Generation @mcp.tool(name="generate_mesh", description="Generate computational mesh using Foam-Agent internals.") def generate_mesh(requirements: str, case_dir: str = "./output", mesh_mode: str = "gmsh", custom_mesh_path: str | None = None): try: from src.config import Config from src.main import initialize_state from src.nodes.meshing_node import meshing_node config = Config() config.case_dir = case_dir state = initialize_state(user_requirement=requirements, config=config, custom_mesh_path=custom_mesh_path) if mesh_mode == "custom": state["mesh_type"] = "custom_mesh" elif mesh_mode == "gmsh": state["mesh_type"] = "gmsh_mesh" else: state["mesh_type"] = "standard_mesh" res = meshing_node(state) return {"success": True, "result": res, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Simulation Generation and Run @mcp.tool(name="generate_simulate", description="Write inputs and run simulation via Foam-Agent graph.") def generate_simulate(requirements: str, case_dir: str = "./output", custom_mesh_path: str | None = None, run_mode: str = "auto"): try: from src.config import Config from src.main import create_foam_agent_graph, initialize_state config = Config() config.case_dir = case_dir state = initialize_state(user_requirement=requirements, config=config, custom_mesh_path=custom_mesh_path) if custom_mesh_path: state["mesh_type"] = "custom_mesh" if run_mode == "local": state["cluster_info"] = None elif run_mode == "hpc": state["cluster_info"] = {"scheduler": "slurm"} workflow = create_foam_agent_graph().compile() workflow.invoke(state) return {"success": True, "result": {"case_dir": config.case_dir}, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
-MCP Tool for Velocity Visualization @mcp.tool(name="visualize_velocity", description="Post-process and visualize velocity (|U|, streamlines, slices).") def visualize_velocity(case_dir: str, plot_type: str = "magnitude", plane: str | None = "xy"): try: from src.config import Config from src.main import initialize_state from src.nodes.visualization_node import visualization_node config = Config() config.case_dir = case_dir state = initialize_state(user_requirement="", config=config, custom_mesh_path=None) state["case_dir"] = case_dir state["visualization_request"] = {"plot_type": plot_type, "plane": plane} vis_res = visualization_node(state) return {"success": True, "result": vis_res, "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
+Appendix D Integration Task Set
+This part documents the task set used in the integration experiments with RepoMaster and OpenAgents.
+The goal of this task set is to evaluate a practical integration question: when Code2MCP supplies MCP services from repositories, can existing agent systems (1) discover the relevant service/tool, (2) invoke it with valid inputs, and (3) complete tool-usage workflows with stable, non-trivial outputs under the same interaction and runtime budgets as the baselines.
+We intentionally keep the task set focused on integration behavior (selection + invocation + composition) rather than using it as a comprehensive benchmark of general agent capability.
+Task sources and construction.
+We construct a fixed set of tasks by first collecting common task types used in prior agent and tool-usage evaluations (e.g., API-oriented code understanding, single-tool execution, and multi-step tool composition), and then instantiating each type into concrete prompts that are executable and checkable in our MCP setting.
+We avoid embedding method-specific hints: task statements do not reference Code2MCP, do not assume any particular tool names, and do not provide privileged repository internals beyond what an agent can obtain through its normal repository discovery interfaces.
+To ensure coverage of both general-purpose and scientific settings, we include a portion of domain-oriented tasks that require using computational libraries via MCP tools, while keeping the task phrasing at the level of user intent rather than narrative case-study descriptions.
+What a task looks like.
+Each task is specified in a uniform template with: (1) a natural-language goal, (2) allowed inputs , (3) constraints relevant to tool usage, and (4) an explicit success condition that is verifiable from tool call traces and returned outputs.
+The success condition is defined at the tool level: a task is successful only if the agent completes the required tool invocations and the final invocation returns a schema-valid JSON object with three fields (success, result, error), where success is true and result is non-empty/non-null and not a direct echo of the input.
+This rule is used to rule out superficial “success” responses that do not reflect meaningful tool execution.
+Task coverage.
+The task set spans three interaction patterns that commonly arise in MCP-based workflows.
+First, tool discovery and selection tasks require the agent to identify which tool corresponds to a stated goal and to infer the minimal valid input format from tool metadata.
+Second, single-step invocation tasks require a single correct tool call that produces a concrete computed output under the environment constraints.
+Third, multi-step composition tasks require two or more tool invocations where later inputs depend on earlier outputs, testing whether intermediate results are correctly propagated and whether the agent can recover from minor argument mismatches by adjusting inputs based on returned error messages.
+We include a small number of “negative” tasks only to probe robustness of invocation behavior; they are evaluated with the same success rule.
+Repository buckets. To analyze the impact of dependency complexity and project style, we group repositories into four buckets based on static metadata:
+  • Low-dependency libraries: at most 10 Python dependencies inferred from requirements.txt or environment.yml, and primarily library-style public APIs.
+  • Medium-dependency scientific libraries: between 11 and 30 Python dependencies with a library-style API and no substantial external system requirements.
+  • High-dependency libraries: more than 30 Python dependencies or additional system-level requirements (e.g., MPI stacks, GPU-accelerated frameworks, or complex external services).
+  • Script-heavy projects: projects whose primary functionality is exposed via CLI scripts, notebooks, or workflow-style pipelines rather than a stable library API; these are identified using repository structure, entry points, and README descriptions.
+Evaluation protocol and fairness controls.
+All methods are evaluated on the same fixed task set with identical task statements and success criteria.
+We do not modify RepoMaster or OpenAgents planning algorithms for individual tasks, and we do not provide additional hints to methods that use Code2MCP-generated tools.
+Each run is executed with the same global budgets (maximum interaction steps and wall-clock time) and with the same runtime restrictions used elsewhere in the paper.
+We log (1) the selected tools, (2) all tool inputs/outputs, (3) per-call durations, and (4) failure reasons (timeouts, schema violations, empty results, or runtime errors).
+When reporting aggregate results, we compute task success rate as the fraction of tasks that meet the success condition within budget, and we additionally report efficiency metrics consistent with the main paper.
+Appendix E Agent Roles and System Prompts
+This section outlines the roles of the specialized agents within the Code2MCP framework.
+Environment Agent This agent rapidly provisions a minimal, isolated runtime for the repository, with minimal dependencies and a short smoke test; if setup fails, propose one lightweight, auditable fallback without modifying the repository; keep all steps reproducible and pragmatic.
+Environment System Prompt - Prefer Conda; use venv only if Conda is unavailable or clearly unsuitable.
+- Detect dependency sources by priority: environment.yml > requirements.txt > pyproject.toml > setup files; never guess hidden dependencies.
+- Pin versions when explicit; otherwise install the minimal viable set.
+Prefer CPU wheels unless GPU is explicitly required.
+- Normalize cross-platform behavior; avoid absolute paths; use relative POSIX-like paths; ensure UTF-8 locale.
+- Smoke test: print Python version and platform; import fastmcp; attempt to import the project’s top-level package or a primary CLI; exit code 0 indicates pass.
+- On failure, capture exact command, exit code, last 80 lines of error logs, and timing; propose exactly one minimal remedy (e.g., switch to venv, install a single missing package, try one version pin, extend timeout once).
+- Apply at most one fallback; never change repository code; do not write outside the workspace; do not weaken security (e.g., no SSL bypass).
+- Cache wheels where possible; avoid global pollution; record reproducible commands and resolved versions.
+- Default to offline validation; if network is strictly required, justify briefly and bound the scope.
+- Emit a compact environment report (platform, Python, manager, explicit deps, resolved pins, pass/fail).
+Code Analysis Expert This agent performs static analysis to shape the repository into a compact, high-value capability surface, selecting stable public functionality, filtering out test/demo code, and producing a concise plan aligned with predefined domains, categories, and solvers.
+Code Analysis System Prompt - Ingest repository signals (structure, import graph, README/docstrings, CLI entry points; DeepWiki if available) to identify stable public APIs suitable as MCP tools.
+- Prefer documented, side-effect-bounded surfaces; exclude tests, internals, notebooks, long demos unless clearly valuable and controllable.
+- Define crisp tool boundaries: explicit inputs/outputs, preconditions/postconditions, resource needs (CPU/GPU/memory/time), and I/O constraints.
+- Note minimal adapter needs (path normalization, dtype coercion, lazy imports) and hazards (network access, file mutation, global state).
+- Summarize fragilities (optional deps, platform quirks) and propose guards (timeouts, argument validation, deterministic seeds).
+- Also produce a case description: case name, case domain, case category, and case solver, using a consistent taxonomy across repositories.
+- Output a compact plan for generation: candidate tools (name, brief description, inputs with types/defaults, outputs, idempotency, side effects) and environment assumptions.
+Keep it actionable and minimal.
+Code Generation Expert This agent synthesizes a robust MCP service from the analysis plan with clean design, consistent interfaces, graceful failure handling, and immediate executability, defining clear tool endpoints, enforcing explicit typed parameters and standardized returns, and avoiding test or example tooling.
+Code Generation System Prompt - Produce clean, runnable Python (no Markdown fences).
+Use FastMCP to build the MCP service.
+- Implement create_app() that returns the service; register tools with concise names and user-facing descriptions.
+- For every tool: explicit, typed parameters (no *args/**kwargs); validate inputs; JSON-serializable outputs.
+- Standard return shape: success: bool, result: any or null, error: string or null.
+- Handle optional dependencies via lazy imports; emit helpful errors without crashing the service; prefer CPU fallbacks when reasonable.
+- Ensure deterministic defaults (fixed seeds when relevant); avoid hidden global state; restrict file I/O to the workspace with existence/size/extension checks.
+- Design for cross-platform paths; avoid shell-specific behavior; bound execution time and memory.
+- Do not generate tests as tools.
+Expose a small set of high-value, composable endpoints; avoid overexposing internals.
+- Add lightweight logging (tool name, argument schema, durations) and minimal version metadata to aid troubleshooting.
+Senior Software Engineer This agent diagnoses failures and applies the smallest auditable change that restores correctness while preserving public contracts, deciding between direct fix and regeneration, using strict complete-file replacement, and avoiding multi-file edits or prose in outputs.
+Review & Auto-Fix System Prompt - Triage failures: import/env, type/contract, path/I-O, dependency/version, timeout/perf, platform.
+- Choose minimal direct fix vs.
+regeneration; provide a one-line rationale and confidence (low/med/high).
+Prefer adapter-boundary mitigations (lazy import, existence checks, parameter coercion).
+- Apply strict complete-file replacement for the single target file; return pure code only; do not alter unrelated sections.
+- Preserve interface contracts and standardized error shapes; add narrow guards instead of broad catches.
+- Enforce cross-platform path handling and deterministic behavior; do not introduce external network calls or new side effects.
+- Optionally add a tiny internal sanity check if it prevents recurrence without bloat.
+- Bound attempts (<= B).
+If still failing, emit a compact escalation note: failing step, last command, error logs tail, and the next best single remediation.
+Final Agent This agent consolidates artifacts and workflow logs into developer-facing documentation and delivery notes that are precise, reproducible, and integration-ready, producing a concise README with installation, quick start, key tools, troubleshooting guidance, and references.
+Final Agent System Prompt - Write a concise developer README (Markdown) including: 1) Overview and value; roles of MCP and FastMCP; supported OS.
+2) Minimal reproducible install (Conda/venv), commands, pinned dependencies, offline notes; Windows PowerShell and Linux shell variants.
+3) Quick start to launch the service and call 2–3 key tools with copy-pasteable snippets and basic error handling.
+4) Tool list: endpoint name, one-line description, key parameters (types/defaults), return shape, idempotency/side effects, typical runtime class.
+5) Troubleshooting: environment/import issues, optional deps, timeouts, path problems, permissions, CPU/GPU enablement; any bounded network caveats.
+6) Reproducibility and telemetry: how to capture environment report, versions, minimal repro commands; where logs/artifacts live.
+7) License and compliance notes: repository license, usage constraints, safety guardrails.
+- Keep structure clear, steps verifiable, and assumptions minimal; prioritize essentials for successful adoption and integration.
+E.1. GPT-4 Template Baseline Prompt
+In the GPT-4 template baseline, we use a single system prompt that instructs one GPT-4 model to generate a complete MCP service implementation for a given repository. The full prompt is given below.
+GPT-4 Template Baseline System Prompt You are a single GPT-4 model acting as an autonomous MCP wrapper generator for a Python code repository.
+You receive: - A short task description.
+- The repository README (possibly truncated).
+- A list of key source files selected by static analysis, each with its file path and content.
+- A minimal description of the MCP protocol and the FastMCP library.
+Your goal is to produce a single, self-contained Python MCP server that exposes a small set of high-value tools from this repository.
+Requirements: 1.
+Design of tools - Identify 3–8 core, stable capabilities that are useful to expose as tools.
+- For each tool, choose a concise, descriptive name and a one-sentence natural-language description.
+- Specify explicit, typed parameters (no *args or **kwargs); include default values where reasonable.
+- Define clear, JSON-serializable return values.
+Avoid returning raw Python objects that cannot be serialized.
+2.
+Implementation with FastMCP - Use the FastMCP library to implement the MCP server.
+- Implement a create_app() function that constructs and returns the MCP application.
+- Register each tool with FastMCP, binding it to the underlying repository functions or classes.
+- Prefer importing from public, documented APIs instead of private or test modules.
+- Use lazy imports for heavy optional dependencies when necessary, and return a helpful error message if a required package is missing.
+3.
+Robustness and safety - Validate all inputs (types, value ranges, file existence) before calling library functions.
+- Handle exceptions inside each tool and return a standardized error object instead of raising uncaught exceptions.
+- The standardized response format for each tool is: { "success": bool, "result": any or null, "error": string or null }.
+- Avoid network access unless the repository explicitly requires it.
+- Restrict file I/O to the workspace; do not write to absolute paths outside the project.
+4.
+Coding style - Write clean, idiomatic Python 3 code with clear function boundaries.
+- Add short docstrings for each tool describing its purpose, parameters, and return value.
+- Do not generate tests, examples, or documentation as part of the MCP server file.
+- Do not include Markdown fences such as ““‘; output only raw Python code.
+5.
+Output format - Return a single Python file containing the complete MCP server implementation.
+- Do not include explanations, comments to the user, or multiple files; only the final server code that can be saved as e.g.
+mcp_service.py.